### v1

# Towards Linear Spanners in All Temporal Cliques

Abstract
Many real-world networks, like transportation networks and social networks, are dynamic in the sense that the edge set may change over time, but these changes are known in advance. This behavior is captured by the temporal graphs model, which has recently become a trending topic in theoretical computer science. A core open problem in the field is to prove the existence of linear-size temporal spanners in temporal cliques, i.e., sparse subgraphs of complete temporal graphs that ensure all-pairs reachability via temporal paths. So far, the best known result is the existence of temporal spanners with $\operatorname{\mathcal{O}}(n\log n)$ many edges. We present significant progress towards proving that linear-size temporal spanners exist in all temporal cliques.
We adapt techniques used in previous works and heavily expand and generalize them to provide a simpler and more intuitive proof of the $\operatorname{\mathcal{O}}(n\log n)$ bound. Moreover, we use our novel approach to show that a large class of temporal cliques, called edge-pivot graphs, admit linear-size temporal spanners. To contrast this, we investigate other classes of temporal cliques that do not belong to the class of edge-pivot graphs. We introduce two such graph classes and we develop novel techniques for establishing the existence of linear temporal spanners in these graph classes as well.
keywords:
temporal graphs, temporal clique, temporal spanner, reachability, graph connectivity, graph sparsification
category:
[…]
1 Introduction
Many real-world networks, like transportation networks with scheduled train connections or social networks with repeating meeting schedules, are dynamic, i.e., their edge-set can change over time. To address this, recent research in theoretical computer science has shifted from static graphs to temporal graphs [15, 18]. In these graphs it is assumed that the edge-set can change in every time step but the schedule of the availability of edges is known in advance. Various problems have been studied in this model, including network redesign [10, 12, 13], vertex cover [1, 11, 14], and influence maximization [9]. A general theme is that insights from the static graph setting are only of limited value in temporal graphs, since the temporal availability of the edges makes these problems much harder or even infeasible to solve. For example, even fundamental theorems like Menger’s Theorem do not hold for temporal graphs [17]. Also, one of the most fundamental problems in graphs, that of finding a sparse spanner, is much harder in temporal graphs. A sparse spanner is a small subset of the edge-set that ensures all-pairs connectivity. In static graphs, it is well-known that linear-size spanners exist and can be computed efficiently, e.g., minimum spanning trees [8].
Finding sparse temporal spanners is much more intricate in temporal graphs, since there paths do not necessarily compose. More specifically, Kempe, Kleinberg, and Kumar [15] showed that there is a class of temporal graphs having $\Omega(n\log n)$ many edges, such that no single edge can be removed while preserving temporal connectivity. In the same work, the authors asked, what is the minimum number of edges $c(n)$, such that any temporal graph with $n$ vertices has a temporal spanner using at most $c(n)$ many edges. Almost 15 years later, this question was answered by Axiotis and Fotakis [2] by providing a class of dense but non-complete temporal graphs having $\Theta(n^{2})$ edges, such that any subgraph preserving temporal connectivity necessarily needs to contain $\Omega(n^{2})$ edges as well.
This negative result raises the question if there even exists a natural class of temporal graphs that can be sparsified without losing temporal connectivity. In particular, the question if every temporal clique has a spanner of linear size has remained open for over 20 years now. In a 2019 breakthrough paper, Casteigts, Peters, and Schoeters [6] proved that temporal cliques always admit a spanner of size $O(n\log n)$. Not much progress has been made since then, with most work on the topic veering off towards the study of temporal spanners in random temporal graphs [7], or focusing on low stretch [3, 4] or a game-theoretic setting [5].
In this paper, we present significant progress on the long-standing open problem by showing that a large class of temporal cliques admits linear-size temporal spanners.
1.1 Our Contribution
In the first part of the paper, we adapt techniques used in previous works and heavily expand and generalize them. In particular, in [6] a temporal spanner in cliques is computed by a reduction from a clique to a bi-clique. We deepen this connection by showing that the existence of a linear spanner in cliques and bi-cliques is equivalent. Furthermore, in the reduction of Casteigts, Peters, and Schoeters [6], the authors exploit the property that the set containing the earliest edge of each vertex is a perfect matching of the bi-clique and that this also holds for the set of latest edges. We show that this property can also be assumed when searching for bi-spanners in bi-cliques directly. This yields a simpler and more intuitive proof, that temporal cliques admit temporal spanners of size $O(n\log n)$.
Moreover, we reconsider the notion of a pivot-vertex [6, 7]. This is a vertex $u$ that can be reached by every vertex in the graph until some time step $t$, and then $u$ can reach every other vertex in the graph. We transfer this notion to bi-cliques, extend it to pivot-edges, and we strengthen it significantly to find small spanners. If a bi-clique contains a pivot-edge, we can iteratively reduce the size of the graph, while maintaining a linear-size temporal spanner, until the instance is solved, or the instance no longer has a pivot-edge. This shows that a large class of graphs admits linear-size temporal spanners.
In the second part of the paper, we investigate temporal cliques that are not edge-pivot graphs. We found two such graph classes, called shifted matching graphs and product graphs, and for both we provide novel techniques that allow us to show that linear-size spanners always exist. Thus we extend the class of temporal cliques that admit linear-size temporal spanners even further. It is open if our classes cover every temporal clique. However, finding other graphs that do not belong to our graph classes seems to be challenging.
1.2 Detailed Discussion of Related Work
The study of temporal spanners was initiated by Kempe, Kleinberg and Kumar [15]. In their seminal paper, they introduced the temporal graphs model and showed that spanners on temporal hypercubes have $\Omega(n\log n)$ edges. Axiotis and Fotakis [2] improved on this by providing a class of graphs with $\Theta(n^{2})$ edges, such that no edge can be removed while maintaining temporal connectivity. Later, Casteigts, Peters, and Schoeters [6] studied temporal cliques and showed that temporal spanners of size $O(n\log n)$ always exist. Their "fireworks" algorithm combines different techniques, such as dismountability and delegation, to reduce the clique to a bipartite graph, and then compute a spanner on that graph, that can then be translated to a spanner of the original graph.
Casteigts, Raskin, Renken and Zamaraev [7] took a different approach. They considered random temporal graphs, where the host graph is a clique, and each edge is added to the temporal graph with probability $p$. They showed a sharp threshold for the generated temporal graph being temporally connected, and, even more surprisingly, they showed that the generated graph will contain an almost optimal spanner of size $2n+o(n)$. On a different note, Bilò, D’Angelo, Gualà, Leucci, and Rossi [4] studied whether temporal cliques admit small temporal spanners with low stretch. Their main result shows that they can always compute a temporal spanner with stretch $O(\log n)$ and $O(n\log^{2}n)$ edges.
2 Preliminaries
For two sets $A,B$, we write ${A}\otimes{B}\coloneqq\{\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}\mid a\in A,b\in B,\text{ and }a\neq b\}$. If $A$ and $B$ are disjoint, we write $A\sqcup B$ for their union. For $n\in\operatorname{\mathbb{N}}^{+}$, let $[n]=\{1,\dots,n\}$.
Let $(V,E)$ be an undirected graph. This together with a labeling function $\operatorname{\lambda}\colon E\to\operatorname{\mathbb{N}}$ forms the temporal graph $G=(V,E,\operatorname{\lambda})$. Semantically, an edge $e\in E$ is present at time $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$. For a vertex-set $U\subseteq V$, write $G[U]$ for the subgraph of $G$ induced by $U$. For $v\in V$, let $N(v)$ denote the set of neighbors of $v$, that is, all vertices $u$ with $\mathopen{}\mathclose{{}\left\{{v},{u}}\right\}\in E$. For $v\in V$ and $u,u^{\prime}\in N(v)$, we define the order ${u}\preceq_{v}{u^{\prime}}$, if the edge $\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}$ is not later than the edge $\mathopen{}\mathclose{{}\left\{{u^{\prime}},{v}}\right\}$, i.e., if $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u}},{{v}}}\right\}}}\right)\leq\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u^{\prime}}},{{v}}}\right\}}}\right)$. We define $\prec_{v}$, $\succeq_{v}$, and $\succ_{v}$ analogously.
[…]
Vertex $u$ can reach vertex $v$ if there is a temporal path from $u$ to $v$. Also, $G$ is called temporally connected, if every vertex can reach every other vertex. A spanner for $G$ is a set $S\subseteq E$, such that $(V,S,{\mathopen{}\mathclose{{}\left.\kern-1.2pt\operatorname{\lambda}\vphantom{\big{|}}}\right|_{S}})$ is temporally connected, where ${\mathopen{}\mathclose{{}\left.\kern-1.2pt\operatorname{\lambda}\vphantom{\big{|}}}\right|_{S}}$ is $\operatorname{\lambda}$ restricted to $S$.
The edges with the smallest and largest label take a special role. For $v\in V$, let $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)\coloneqq\operatorname*{arg\,min}_{u\in N(v)}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{u}}}\right\}}}\right)$ be $v$’s earliest neighbor and let $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({v}}\right)\coloneqq\operatorname*{arg\,max}_{u\in N(v)}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{u}}}\right\}}}\right)$ be $v$’s latest neighbor. We do not consider graphs with isolated vertices and we justify in Section 3 that we only use injective labelings, so this is well-defined. Denote the set of all earliest edges as $\operatorname{\pi^{\text{--}}}\coloneqq\{\mathopen{}\mathclose{{}\left\{{v},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)}}\right\}\mid v\in V\}$ and the set of latest edges as $\operatorname{\pi^{+}}\coloneqq\{\mathopen{}\mathclose{{}\left\{{v},{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({v}}\right)}}\right\}\mid v\in V\}$. If both sets $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ form a perfect matching for $V$, then $G$ is extremally matched. Also, for $v,w\in V$, and $e=\mathopen{}\mathclose{{}\left\{{v},{w}}\right\}\in E$, we write $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)\in[{n}]$ for the index of $e$, when ordering all edges incident to $v$ ascendingly by their label (e.g., if $w=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)$, then $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)=1$).
Let $v\in V$ be a vertex and $t\in\operatorname{\mathbb{N}}$ be a time step. We define the set $\operatorname{In}_{v}^{t}\subseteq V$ to be the vertices reaching $v$ via a temporal path with every label being at most $t$. Respectively, we define the set $\operatorname{Out}_{v}^{t}\subseteq V$ to be the vertices that $v$ can reach via any temporal path starting at or after $t$. As a shorthand, for an edge $e=\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}\in E$, we define $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)=\operatorname{In}_{u}^{\operatorname{\lambda}(e)}\cup\operatorname{In}_{v}^{\operatorname{\lambda}(e)}$ and $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)=\operatorname{Out}_{u}^{\operatorname{\lambda}(e)}\cup\operatorname{Out}_{v}^{\operatorname{\lambda}(e)}$. Note, that we always have $u,v\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$.
We focus only on temporal graphs where the underlying graph is a clique or a complete bipartite graph. In these cases, the set of edges is self-evident and for cliques we simply omit it and write $C=(V,\operatorname{\lambda})$. We, similarly, denote complete bipartite temporal graphs by a triple $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $A$ and $B$ are the two partitions and $\operatorname{\lambda}\colon{A}\otimes{B}\to\operatorname{\mathbb{N}}$ is the corresponding labeling function. When necessary, we refer to the set of edges as $E(C)\coloneqq{V}\otimes{V}$ or $E(D)\coloneqq{A}\otimes{B}$. We call $C$ a temporal clique and $D$ a temporal bi-clique. Finally, we define a directed analogue of spanners for temporal bi-cliques. We call a set $S\subseteq{A}\otimes{B}$ a bi-spanner for $D$, if every $a\in A$ can reach every $b\in B$ through a temporal path that is contained in $S$.
For $e=\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}$ in temporal cliques or bi-cliques, notice that all vertices must be in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ or $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$, since all vertices have an edge to $u$ or $v$, which puts this vertex in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ or $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$, depending on the time-label. See Figure 0(a) for a visualization of $\operatorname{In}$ and $\operatorname{Out}$.
(a) Blue edges and paths have labels earlier than $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$, red ones are later.
(b) A bi-clique where $a_{1}$ cannot reach $a_{2}$ after time $2$ or can be reached by $a_{2}$ until time $4$, so $\operatorname{In}_{v}^{3}\cup\operatorname{Out}_{v}^{3}\neq V$.
Figure 1: Examples for the sets $\operatorname{In}$ and $\operatorname{Out}$.
{observation}
Let $G=(V,E,\operatorname{\lambda})$ be a temporal clique or bi-clique. For all $e\in E$, we have $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)=V$.
Note that for the case of bi-cliques considering both endpoints of $e$ is crucial to ensure that every vertex has a direct edge to an endpoint. Otherwise, we simple examples exist having no vertex $v$ with time $t$ with $\operatorname{In}_{v}^{t}\cup\operatorname{Out}_{v}^{t}=A\sqcup B$. See Figure 0(b).
Remark 2.1.
For $v\in A\sqcup B$ and $t\in\operatorname{\mathbb{N}}$, the analogue to section 2 (i.e., $\operatorname{In}_{v}^{t}\cup\operatorname{Out}_{v}^{t}=A\sqcup B$) does not necessarily hold, as one can see in Figure 0(b).
Finally, for $n\in\operatorname{\mathbb{N}}$, let $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that every temporal clique with $n$ vertices admits a spanner with at most $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges. Analogously, let $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that any bi-clique in which both sides have $n$ vertices admits a bi-spanner with at most $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges. We study the open problem whether $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ holds.
3 Cliques and Bi-Cliques
In this section, we prove that the asymptotic behavior of spanners in cliques differs from the asymptotic behavior of bi-spanners in bi-cliques only by a constant factor (i.e., we prove $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\Theta(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right))$). First, we show that $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right))$. After this, we show $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\Omega(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right))$ by proving each inequality of $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({\frac{n}{2}}}\right)\geq\frac{1}{4}\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ separately. Bi-cliques and bi-spanners were already discussed in [6], when deriving $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n\log n)$. They called these instances residual as they remained once their other techniques were not applicable anymore. Our result can be seen as evidence that the connection between cliques and bi-cliques is fundamental, beyond being a step in their algorithm. First, we state the main result of this section.
Theorem 3.1.
The worst-case minimum size of a spanner for temporal cliques and temporal bi-cliques only differs by a constant factor, that is, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\Theta(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right))$.
We start with a technical lemma that allows us to focus on injective labeling functions for the rest of the paper. Note that the statement holds for all temporal graphs.
Lemma 3.2.
Let $G=(V,E,\lambda)$ be a temporal graph. Then there is an injective labeling function $\lambda^{\prime}$ s.t. every spanner $S\subseteq E$ for $(V,E,\lambda^{\prime})$ is also a a spanner for the original graph $G$. In case the original graph $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ is a bi-clique, the same holds for a bi-spanner $S$.
Proof 3.3.
Assume $\lambda$ is not injective. To fix that, we simply arbitrarily order edges with the same label. For a simple construction, assume $f\colon E\to[|E|]$ is an arbitrary numbering of the edges and for all $e\in E$ set $\lambda^{\prime}(e)\coloneqq\lambda(e)\cdot(|E|+1)+f(e)$. To see that $\lambda^{\prime}$ is injective, first take distinct $e_{1},e_{2}\in E$ with $\lambda(e_{1})=\lambda(e_{2})$. In this case, $\lambda^{\prime}(e_{1})-\lambda^{\prime}(e_{2})=(\lambda(e_{1})+f(e_{1}))-(\lambda(e_{2})+f(e_{2}))=f(e_{1})-f(e_{2})$ and since $f$ is injective this difference is not 0, thus we have $\lambda^{\prime}(e_{1})\neq\lambda^{\prime}(e_{2})$. Now assume we have $e_{1},e_{2}\in E$ with $\lambda(e_{1})<\lambda(e_{2})$. Then we have
$\lambda^{\prime}(e_{1})$ $=\lambda(e_{1})\cdot(|E|+1)+f(e)$
$\leq\lambda(e_{1})\cdot(|E|+1)+|E|$
$<(\lambda(e_{1})+1)\cdot(|E|+1)$
$\leq\lambda^{\prime}(e_{2}).$
Also observe that for all $e_{1},e_{2}\in E$ we have $\lambda^{\prime}(e_{1})\leq\lambda^{\prime}(e_{2})$ then $\lambda(e_{1})\leq\lambda(e_{2})$. Thus if $P$ is a valid temporal path w.r.t. $(V,E,\lambda^{\prime})$, then $P$ is also a valid temporal path w.r.t. $G$. Conclude that if $S$ is a spanner in $(V,E,\lambda^{\prime})$, then $S$ is a spanner in $G$.
Notice that the same construction works for bi-spanners, since we can again break ties arbitrarily and all remaining temporal paths existed in the original instance.
We can now relate minimum spanners in cliques to minimum bi-spanners in bi-cliques.
Lemma 3.4.
Let $C=(V,\operatorname{\lambda})$ be a temporal clique. There is a temporal bi-clique $D=(A,B,\operatorname{\lambda}^{\prime})$ with $|V|=|A|=|B|$ s.t. for any bi-spanner $S_{D}$ in $D$ there is a spanner $S_{C}$ in $C$ with $|S_{C}|\leq|S_{D}|$.
Proof 3.5.
Remember, by Lemma 3.2, we can assume that $\operatorname{\lambda}$ is an injection.
For every vertex $v\in V$, we create a duplicate vertex $v^{\prime}$. Call the set of all duplicates $V^{\prime}$. We define the edge label between two vertices as the label between the original vertices. For the edge label between a vertex in $V$ and its copy, we choose a label higher than all in the original graph. That is, to define $\operatorname{\lambda}^{\prime}$, choose $\mu>\max_{e\in{V}\otimes{V}}\operatorname{\lambda}(e)$ and set for every $u,v\in V$
$\operatorname{\lambda}^{\prime}(\mathopen{}\mathclose{{}\left\{{u},{v^{\prime}}}\right\})\coloneqq\begin{cases}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u}},{{v}}}\right\}}}\right),&\text{if}u\neq v;\\ \mu,&\text{if }u=v.\end{cases}$
An example of this transformation can be seen in Figure 2.
(a) A temporal clique $C$.
(b) The corresponding temporal bi-clique $D$.
Figure 2: Example of a temporal clique $C$ transformed into a temporal bi-clique $D$.
Consider the bi-clique $D=\mathopen{}\mathclose{{}\left({V},{V^{\prime}},{\operatorname{\lambda}^{\prime}}}\right)$ and assume $S_{D}$ is a bi-spanner in $D$. Let $S_{C}\coloneqq\operatorname{\lambda}^{-1}(\operatorname{\lambda}^{\prime}(S_{D})\setminus\{\mu\})$, that is, all edges in $C$ for which an edge with the same time-label is included in $S_{D}$. As $\operatorname{\lambda}$ is injective, we directly observe that $|S_{C}|\leq|S_{D}|$.
To show that $S_{C}$ is a spanner in $C$, consider $u,v\in V$. We need to show, that there is a temporal $u$$v$-path only using edges of $S_{C}$. As $S_{D}$ is a temporal bi-spanner in $D$, there is a temporal $uv^{\prime}$-path $P\subseteq S_{D}$. Because $\mu$ is the largest label and each vertex has exactly one adjacent edge with label $\mu$, if $P$ contains an edge with time-label $\mu$, it has to end with this edge. If that is the case, consider $P$ without this last edge, which by construction is an $uv$-path in $D$. Since $\lambda$ is injective, notice that the path using the same labels in the same order is a temporal path between $u$ and $v$ in $S_{C}$.
This is enough to arrive at the first half of the tight bound we aim to prove in the section.
Theorem 3.6.
For $n\in\operatorname{\mathbb{N}}$, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
Proof 3.7.
Let $C=(V,\operatorname{\lambda})$ be a temporal clique with $|V|=n$, such that any minimum spanner $S^{*}$ in $C$ obeys $\mathopen{}\mathclose{{}\left|S^{*}}\right|=\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$.
By Lemma 3.4, there is a temporal bi-clique $D$ with $n$ vertices per side such that for any minimum bi-spanner $S_{D}$ there is a spanner $S_{C}$ in $C$ with $|S_{C}|\leq|S_{D}|\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. By definition of $S^{*}$, we conclude $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)=|S^{*}|\leq|S_{C}|\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
We now show that $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. To this end, we take a bi-clique and add edges within the partitions that yield no benefit in the construction of a spanner.
Lemma 3.8.
Let $D=(A,B,\operatorname{\lambda})$ be a temporal bi-clique with $\mathopen{}\mathclose{{}\left|A}\right|=\mathopen{}\mathclose{{}\left|B}\right|$. Then there is a temporal clique $C=(A\sqcup B,\operatorname{\lambda}^{\prime})$ such that for any spanner $S$ in $C$ the set $S\cap({A}\otimes{B})$ is a bi-spanner in $D$.
Proof 3.9.
We define $\operatorname{\lambda}^{\prime}$ such that for distinct $v,w\in A\sqcup B$, we have
$\operatorname{\lambda}^{\prime}(\mathopen{}\mathclose{{}\left\{{v},{w}}\right\})=\begin{cases}0,&\text{if }v,w\in B;\\ 1+\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{w}}}\right\}}}\right),&\text{if }v\in A,w\in B\text{ or }v\in B,w\in A;\\ \mu&\text{if }v,w\in A,\end{cases}$
where $\mu>1+\max_{e\in{A}\otimes{B}}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$. Intuitively, all early edges are between vertices in $B$ and all late edges are between vertices in $A$. The original edges keep their ordering.
Let $S$ be a spanner of $C$. To show $S\cap({A}\otimes{B})$ is a bi-spanner in $D$, let $a\in A$ and $b\in B$. As $S$ is a spanner, there is a temporal $ab$-path $P\subseteq S$. If $P$ contains any edges within $A$, no edge to $B$ can be used afterward, due to the labeling. Similarly, once the path arrives at $B$, it is already too late to use any edge within $B$. Thus, $P\subseteq{A}\otimes{B}$ and $S\cap({A}\otimes{B})$ is a bi-spanner in $D$.
This yields the following relationship between clique and bi-clique spanner sizes.
Theorem 3.10.
For $n\in\operatorname{\mathbb{N}}$, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
Proof 3.11.
Let $D=(A,B,\operatorname{\lambda})$ be a temporal bi-clique with $|A|=|B|=n$, such that any minimum bi-spanner $S^{*}$ in $D$ obeys $\mathopen{}\mathclose{{}\left|S^{*}}\right|=\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. By Lemma 3.8, there is a temporal clique $C=(A\sqcup B,\operatorname{\lambda}^{\prime})$ such that any spanner $S$ in $C$ satisfies $\mathopen{}\mathclose{{}\left|S}\right|\geq\mathopen{}\mathclose{{}\left|S^{*}}\right|$. Since the vertex set of $C$ has size $2n$, there is a spanner of size $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)$ in $C$, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
To conclude this section, we just need to show, that $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\frac{1}{4}\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({2n}}\right)$. To that end, we show the slightly more general statement that for any $k\in\operatorname{\mathbb{N}}^{+}$, we have $k^{2}\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({kn}}\right)$, which we achieve by splitting the instances into smaller subinstances.
Lemma 3.12.
For all $k,n\in\operatorname{\mathbb{N}}^{+}$, we have $k^{2}\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({kn}}\right)$.
Proof 3.13.
Consider any temporal bi-clique $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $\mathopen{}\mathclose{{}\left|A}\right|=\mathopen{}\mathclose{{}\left|B}\right|=kn$. Consider a partition of $A$ and $B$ into $k$ equally sized blocks $A_{1},A_{2},\dots,A_{k}\subseteq A$ and $B_{1},B_{2},\dots,B_{k}\subseteq B$. For any $i,j\in[k]$ define the temporal graph $D_{i,j}=D[A_{i}\sqcup B_{j}]$ to be the sub-graph of $D$ induced on the respective blocks with the labeling $\operatorname{\lambda}$ restricted appropriately and let $S_{i,j}$ be a minimum bi-spanner on $D_{i,j}$. We now set $S\coloneqq\bigcup_{i,j\in[k]}S_{i,j}$.
For any $i,j\in[k]$, we have $\mathopen{}\mathclose{{}\left|A_{i}}\right|=\mathopen{}\mathclose{{}\left|B_{j}}\right|=n$ and therefore $\mathopen{}\mathclose{{}\left|S_{i,j}}\right|\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ and $S$ has appropriate size (i.e., $\mathopen{}\mathclose{{}\left|S}\right|\leq k^{2}\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$). Let $i,j\in[k]$ and $a\in A_{i},b\in B_{j}$. Then there is a temporal path from $a$ to $b$ using only edges of $S_{i,j}\subseteq S$, so $S$ is a bi-spanner for $D$.
Remark 3.14.
We later strengthen this statement using tools developed in section 4. Concretely, considering $D_{i}\coloneqq D[A_{i}\cup B]$, we show $2k(k-1)n+k\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({kn}}\right)$.
To plug all pieces together, we need a final property about $\operatorname{\mathcal{D}}$.
Lemma 3.15.
The worst-case size of spanners for bi-cliques is monotonic in the size of the bi-clique, that is for all $n\in\operatorname{\mathbb{N}}$, we have $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n+1}}\right)$. The same holds for $\operatorname{\mathcal{C}}$.
Proof 3.16.
We only prove the statement for bi-cliques and $\operatorname{\mathcal{D}}$, the proof for cliques and $\operatorname{\mathcal{C}}$ is analogous. Let $n\in\operatorname{\mathbb{N}}$ and let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique of size $n$. To construct a bi-clique of size $n+1$, we clone one vertex on each side. To be exact, take $a\in A$ and $b\in B$ and let
$A^{\prime}$ $\coloneqq A\cup\{a^{\prime}\},$
$B^{\prime}$ $\coloneqq B\cup\{b^{\prime}\},$
[…]
$D^{\prime}$ $\coloneqq(A^{\prime}\sqcup B^{\prime},\operatorname{\lambda}^{\prime}).$
Now any spanner $S^{\prime}$ for $D^{\prime}$ can be transformed into a spanner $S$ of smaller or equal size for $D$, by replacing any edge in $S^{\prime}$ adjacent to $a^{\prime}$ or $b^{\prime}$ with the corresponding edge adjacent to $a$ respective $b$. Any temporal path in $S^{\prime}$ now yields a temporal path in $S$ where we can use $a$ and $b$ instead of $a^{\prime}$ and $b^{\prime}$ since the labels of the original and cloned edges are the same.
Now all pieces are in place to puzzle together the proof of our main theorem.
Proof 3.17 (Proof of Theorem 3.1).
Let $n\in\operatorname{\mathbb{N}}$ and $c\in\{0,1,2,3\}$ be such that $n\geq 3$ and $n+c$ is a multiple of 4. In the following equation, we use Lemma 3.15 in the first and last step, apply Lemma 3.12 in the third step and Theorem 3.10 in the forth step, to obtain
$\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n+c}}\right)=\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({4\cdot\frac{n+c}{4}}}\right)\leq 16\cdot\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({\frac{n+c}{4}}}\right)\leq 16\cdot\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({\frac{n+c}{2}}}\right)\leq 16\cdot\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right).$
Theorem 3.6 directly gives us $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$, allowing us to conclude $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\Theta(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right))$.
4 Structure of Bi-Clique Instances
Using Theorem 3.1, from now on we only consider bi-cliques. Our first structural insight for this type of instance was already used by Casteigts, Peters, and Schoeters [6] on special reduced instances of bi-cliques. Remember, that we call an instance extremally matched if the set of all earliest edges $\operatorname{\pi^{\text{--}}}$ and the set of all latest edges $\operatorname{\pi^{+}}$ each form a perfect matching. We give a general reduction rule which reduces the graph in case the instance is not extremally matched. This structural property, that we can assume the existence of such a perfect matching or reduce the instance, is useful for almost all further constructions. For example, it enables us to reduce the instance size, when the number of vertices per side is not equal. Additionally, we can think of the sets $A$ and $B$ of the bi-clique as interchangeable. For example, consider any set of edges that includes temporal paths from some set $A^{\prime}\subseteq A$ to all vertices in $A$. We can extend these paths to reach all vertices in $B$ by including the set of all latest edges, because it is a perfect matching between $A$ and $B$.
We start with proving that every $b\in B$ is incident to at most one edge in $\operatorname{\pi^{\text{--}}}$ or we can reduce. The same holds for any $a\in A$ with respect to $\operatorname{\pi^{+}}$. For our reduction rules, we adapt the concept of dismountability on temporal cliques, first defined in [6], to temporal bi-cliques.
Definition 4.1 (dismountable).
Vertex $a\in A$ is dismountable if there is $a^{\prime}\in A$ with $a\neq a^{\prime}$, s.t. ${a}\preceq_{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}{a^{\prime}}$. Vertex $b\in B$ is dismountable if there is $b^{\prime}\in B$ with $b\neq b^{\prime}$, s.t. ${b^{\prime}}\preceq_{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}{b}$.
Intuitively, vertex $a$ can delegate its obligation to reach all vertices in $B$ to vertex $a^{\prime}$ in exchange for including two edges in the bi-spanner. Similarly, vertex $b$ can be reached by all vertices in $A$ that can reach vertex $b^{\prime}$ if we include the two necessary edges in the bi-spanner.
Note that this corresponds to the definition in [6], but transferred to the case of bi-spanners, where all vertices in $A$ only need to reach vertices in $B$. Concretely, this means that vertex $a$ can reach vertex $a^{\prime}$ via the earliest edge of $a^{\prime}$ on a temporal path of length 2. Therefore, if this configuration occurs in the bi-clique, we can remove $a$ and include the edges $\mathopen{}\mathclose{{}\left\{{a},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}}\right\}$ and $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}}\right\}$ into our bi-spanner. In the final solution for the whole bi-clique, vertex $a$ then has a temporal path to every vertex $b\in B$ via vertex $a^{\prime}$. Similarly, if a vertex $b\in B$ is dismountable, every temporal path that ends in a different vertex can be extended to be a temporal path ending in vertex $b$, and we can remove $b$ from the bi-clique.
The following lemma is similar to the one in [6]. For completeness, we prove it again.
Lemma 4.2.
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique. Let $a\in A$, $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a}}\right)=b$, then $a=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)$ or $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)$ is dismountable. Let $b\in B$, $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b}}\right)=a$, then $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({a}}\right)=b$ or $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({a}}\right)$ is dismountable.
Proof 4.3.
Let $a\in A$ with $b=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a}}\right)$. By definition of $\operatorname{\pi^{\text{--}}}$, we have ${\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)}\preceq_{b}{a}$. If $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)\neq a$, $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)$ is dismountable through $a$. The other case follows in the same fashion.
We can apply Lemma 4.2 to show that we can always find a dismountable vertex whenever the number of vertices per side differs.
Lemma 4.4.
For $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, if $|A|\neq|B|$ then a dismountable vertex exists.
Proof 4.5.
If $|A|>|B|$, by the pigeonhole principle, there must be at least two vertices in $A$ with the same earliest neighbor. Similarly, if $|A|<|B|$, there must be at least two vertices in $B$ with the same latest neighbor. Due to Lemma 4.2 there must be a dismountable vertex in these cases.
Since from now on, both partitions of $D$ have the same size after removing dismountable vertices, we write $n\coloneqq|A|=|B|$ and call $n$ the size of the instance. This justifies also that the definition of $\operatorname{\mathcal{D}}$ assumes equal partition sizes and gives the following corollary.
Corollary 4.6.
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ and $|A|\leq|B|$. Then, there is a bi-spanner for $D$ of size $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({|A|}}\right)+2(|B|-|A|)$. Similarly, if $|B|\leq|A|$ there is a bi-spanner of size $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({|B|}}\right)+2(|A|-|B|)$.
Finally, combining our lemmas yields that $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ are extremally matched.
Theorem 4.7 (Extremal Matching).
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique. Then $D$ is extremally matched or there is a dismountable vertex.
Proof 4.8.
We only prove the statement for $\operatorname{\pi^{\text{--}}}$, the proof for $\operatorname{\pi^{+}}$ follows analogously. Suppose there is no dismountable vertex. Let $b\in B$. By definition, $b$ has at least one incident edge in $\operatorname{\pi^{\text{--}}}$, namely $\mathopen{}\mathclose{{}\left\{{b},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)}}\right\}$. If there is an $a\in A$ with $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a}}\right)=b$, due to Lemma 4.2, we have $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({b}}\right)=a$, thus $\pi^{-}$ has at most one edge adjacent to $b$. Therefore $b$ has exactly one incident edge and with Lemma 4.4, $\operatorname{\pi^{\text{--}}}$ is a perfect matching.
After removing dismountable vertices, we now know that the remaining graph must be extremally matched. Since reducing an instance to be extremally matched only requires the inclusion of linearly many edges, we focus from now on extremally matched bi-cliques. This also allows us to switch the roles of $A$ and $B$, which will be useful later on.
Lemma 4.9.
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be an extremally matched bi-clique and let $S$ be a bi-spanner for $D$. If $\operatorname{\pi^{\text{--}}},\operatorname{\pi^{+}}\subseteq S$, then $S$ is also a bi-spanner for $\mathopen{}\mathclose{{}\left({B},{A},{\operatorname{\lambda}}}\right)$.
Proof 4.10.
[…]
With Corollary 4.6 we prove Lemma 4.11. This gives a simpler proof that $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n\log n)$.
Lemma 4.11.
For all $k,n\in\operatorname{\mathbb{N}}^{+}$, we have $2k(k-1)n+k\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({kn}}\right)$.
Proof 4.12.
The proof is similar to the proof of Lemma 3.12. First, consider any temporal bi-clique $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $|A|=|B|=kn$. We partition $A$ into equally sized blocks $A_{1},A_{2},\ldots,A_{k}\subseteq A$. For $i\in[k]$, let $D_{i}\coloneqq D[A_{i}\sqcup B]$ with appropriately restricted labeling.
For every $D_{i}$ consider a minimum spanner $S_{i}$, which satisfies $|S_{i}|\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)+2(k-1)n$ by Corollary 4.6. Set $S\coloneqq\bigcup_{i\in[k]}S_{i}$. Then, $S$ has the correct size and for every $i\in[k],a\in A_{i}$, and $b\in B$, there is a temporal $ab$-path using only edges of $S_{i}\subseteq S$.
def biSpanner($D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$):       if $|A|=1$ then  return $E(D)$ partition $A$ into non-empty $A_{1},A_{2}$ with $|A_{1}|=\lfloor|A|/2\rfloor$ $D_{1}\leftarrow D[A_{1}\sqcup B]$, $D_{2}\leftarrow D[A_{2}\sqcup B]$ $(D_{1}^{*},S_{1}^{*})\leftarrow$ dismount($D_{1}$), $(D_{2}^{*},S_{2}^{*})\leftarrow$ dismount($D_{2}$) return biSpanner($D_{1}^{*}$) $\cup$ biSpanner($D_{2}^{*}$) $\cup$ $S_{1}^{*}$cup $S_{2}^{*}$
Figure 3: Our algorithm for computing $\operatorname{\mathcal{O}}(n\log n)$ bi-spanners based on Lemma 4.11. The function dismount takes a bi-clique, exhaustively dismounts all dismountable vertices, and returns the remaining instance and the included edges. By Lemma 4.4, this dismounts at least until $|A|=|B|$.
If we now choose $k>1$, we get $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n\log n)$ using the master theorem [8]. Note that we can also implement all the necessary reductions to find such a spanner in linear time in the number of edges in the temporal clique. This achieves the same spanner size as in [6] but with a simpler algorithm and proof. See Figure 3 for an outline of the algorithm.
Corollary 4.13.
We have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n\log n)$ and there is an algorithm that computes a spanner of size $\operatorname{\mathcal{O}}(n\log n)$ in time $\operatorname{\mathcal{O}}(n^{2})$.
5 Pivotable Bi-Cliques
An important structure that was used extensively in [6, 7] is the notion of a pivot. A pivot is a vertex that can be reached by all vertices until a certain time point $t$ and that can also reach all other vertices starting at time point $t$. In [7] it is shown that this structure, though intuitively quite restrictive, appears with high probability in random temporal graphs.
However, there are instances without this structure [6]. Thus, we study an even more widely applicable structure, based on a what we call a pivot-edge. We exploit pivot-edges to apply the divide-and-conquer paradigm to find linear-size temporal spanners. To the best of our knowledge, this paradigm has not yet been applied to this problem. As we see later, the absence of a pivot-edge yields strong structural properties and it heavily restricts the class of temporal cliques where no linear spanner can be found using current techniques.
5.1 Pivot-Edges
Recall the definition of $\operatorname{In}$ and $\operatorname{Out}$ in Section 2. We observe that in extremally matched bi-cliques, the sets $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ and $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ are always distributed evenly between $A$ and $B$.
Lemma 5.1.
Let $G=(A,B,\lambda)$ be an extremally matched bi-clique. For all $e\in{A}\otimes{B}$, we have $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap A}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap B}\right|$ and $\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)\cap A}\right|=\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)\cap B}\right|$.
Proof 5.2.
Consider any $e\in A\otimes B$. Notice that for all $v\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$, we have $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ and $\operatorname{\pi^{\text{--}}}$ is a bijection between $A$ and $B$. So, there is a bijection between $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap A$ and $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap B$. Thus, $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap A}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap B}\right|$. The claim regarding $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ follows analogously.
Next, we show that we can reach all vertices in $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ from $e$ starting at or after $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$, using exactly $\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-1$ edges. The analogous statement holds for the vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$, in the sense that those can reach $e$ at or after $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$, using only $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-1$ many edges.
Lemma 5.3.
Let $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}\in{A}\otimes{B}$. All vertices in $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ can be organized in a tree $T$ rooted at $a$, s.t. $T$ only has edges with label at least $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$ and for each $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ there are temporal paths $a\leadsto v$ and $b\leadsto v$ in $G[T]$. The analogous result holds for vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$.
The proof of Lemma 5.3 is already known [19] and it relies on constructing a foremost tree for $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ (and its inverted version for $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right))$. It was considered in [16] and lately has been used in [7], where foremost trees are applied to temporal cliques.
We can directly use this idea to temporally connect all $u\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ to all $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using a linear number of edges.
Corollary 5.4.
Given an edge $e\in{A}\otimes{B}$, we can connect the vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ to those in $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ by using at most $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3$ many edges.
Proof 5.5.
Since we can organize the vertices $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$, $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ as a tree by Lemma 5.3, all vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ can reach $e$ and, afterwards, can reach $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using the respective trees. Both trees consist of $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-1,\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-1$ many edges respectively. But since $a$ and $b$ appear in both sets being connected by $e$, we double count this edge resulting in $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3$ many edges.
To apply divide-and-conquer, we choose an edge $e\in{A}\otimes{B}$ and include the edges according to Corollary 5.4. Then, we create two sub-instances to connect the remaining vertices.
Theorem 5.6.
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be an extremally matched bi-clique with $n\coloneqq\mathopen{}\mathclose{{}\left|A}\right|=\mathopen{}\mathclose{{}\left|B}\right|$. There exists a bi-spanner in $D$ with size at most
$\min_{e\in{A}\otimes{B}}\mathopen{}\mathclose{{}\left(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n-\frac{\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|}{2}}}\right)+\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n-\frac{\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|}{2}}}\right)+2\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+2\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3}\right).$
Proof 5.7.
Consider any $e\in{A}\otimes{B}$. By Corollary 5.4, there is a set of edges $F_{e}$ such that $|F_{e}|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3$ and all vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ can reach all vertices of $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using the edges in $F_{e}$. To connect the remaining vertices, we create two sub-instances $G^{\prime}_{e}\coloneqq G[(A\setminus\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right))\sqcup B]$ and $G^{\prime\prime}_{e}\coloneqq G[A\sqcup(B\setminus\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right))]$. Let $S^{\prime}_{e}$ be a minimum bi-spanner of $G^{\prime}_{e}$, and $S^{\prime\prime}_{e}$ be a minimum bi-spanner of $G^{\prime\prime}_{e}$. By Lemma 5.1, the smaller sides of $G^{\prime}_{e}$ and $G^{\prime\prime}_{e}$ have size $n-\frac{\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|}{2}$ and $n-\frac{\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|}{2}$, respectively. By Corollary 4.6, we get that $|S^{\prime}_{e}|\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n-\frac{\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|}{2}}}\right)+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|$ and $|S^{\prime\prime}_{e}|\leq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n-\frac{\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|}{2}}}\right)+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|$.
We claim that the set $S_{e}\coloneqq F_{e}\cup S^{\prime}_{e}\cup S^{\prime\prime}_{e}$ is a bi-spanner for $G$. Consider any vertex $a\in A$ and vertex $b\in B$. If vertex $a\notin\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$, then vertex $a$ can reach vertex $b$ via $S^{\prime}_{e}$. If $b\notin\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$, vertex $a$ can reach vertex $b$ via $S^{\prime\prime}_{e}$. Otherwise, $a\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ and $b\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ and so $a$ can reach $b$ via $F_{e}$. Choosing the edge where the set $S_{e}$ is smallest, yields the result.
As $32\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+2\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|-3\leq 8n-$ 3, we see that only a linear number of edges is included to split the instance. With this, we can show that if for some edge $e\in{A}\otimes{B}$ the sets $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ and $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ overlap, we can reduce the instance significantly. We name the set ${\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}$ the pivot-set of $e$. If $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$ for $c\in(0,1]$, we call $e$ a $c$-pivot-edge.
Theorem 5.8 (Pivot Edge).
Let $c\in(0,1]$. Let $D=(A,B,\lambda)$ be an extremally matched bi-clique with $n\coloneqq|A|=|B|$. If there is a $c$-pivot-edge in $D$, we can split the instance into two sub-instances with overall number of vertices reduced by $cn$. For each removed vertex, we include at most $\frac{8}{c}$ edges. If applicable recursively, we get a spanner of size $\frac{16}{c}n$.
Proof 5.9.
Let $e\in{A}\otimes{B}$ be such that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$. Note that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq\mathopen{}\mathclose{{}\left(2+c}\right)n$. By Theorem 5.6, the sub-instances have sides of overall size at most $2n-\frac{2+c}{2}n=\mathopen{}\mathclose{{}\left(1-\frac{c}{2}}\right)n$. This means that $cn$ vertices have been removed using at most $38n-$ 3 edges overall. For each such vertex, we include $\frac{8n-3}{cn}\leq\frac{8}{c}$ edges.
To recursively apply Theorem 5.6, observe that we can reduce any dismountable vertex until both sub-instances are extremally matched. For each vertex removed in this way, we include $2\leq\frac{8}{c}$ edges. As the number of edges included per removed vertex in both ways does not depend on $n$, we include at most $\frac{8}{c}\mathopen{}\mathclose{{}\left|A\sqcup B}\right|=\frac{16}{c}n$ edges.
This theorem tells us that if we repeatedly find such a $c$-pivot-edge $e\in{A}\otimes{B}$ with $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|$ being at least a (previously fixed) $c$-fraction of the current size of the instance, we get a linear-size bi-spanner for our graph. Therefore, if we search for counterexamples for linear-size spanners in temporal bi-cliques, we can assume that for any fixed $c$ no such $c$-pivot-edge is present, that is, for all $e\in{A}\otimes{B}$, we have $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\in\operatorname{o}(n)$.
The following subsection gives an intuition for how restrictive this condition actually is.
5.2 Graph Properties leading to Pivot-Edges
Pivot-edges generalize pivotable vertices and yield a number of new, interesting structural properties in residual instances. In this section, we examine structures that, if present, imply that there is at least one pivot-edge.
We first look at one particular case of pivot-edges, which we call steep edges. Intuitively, an edge is steep, if it belongs to the later edges for one of the incident vertices and to the earlier edges for the other one.
Definition 5.10 ($c$-steep).
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique. For a constant $c\in(0,1]$, we call $e$ a $c$-steep edge, if $\mathopen{}\mathclose{{}\left|\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)-\sigma_{w}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$.
We now show that we can assume edges not to be steep even in artificial instances, by proving that as long as there is a steep edge, we can apply Theorem 5.8.
Lemma 5.11.
Let $c\in(0,1]$. If there is a $c$-steep edge $e\in{A}\otimes{B}$, then $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$.
Proof 5.12.
[…]
$\mathopen{}\mathclose{{}\left|A^{-}}\right|+\mathopen{}\mathclose{{}\left|B^{+}}\right|$ $\leq\frac{1}{2}(\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|)$
$=\frac{1}{2}(\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|)$
$=n+\frac{1}{2}(\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|).$
We now bound $\mathopen{}\mathclose{{}\left|A^{-}}\right|+\mathopen{}\mathclose{{}\left|B^{+}}\right|$ from below. We have $\mathopen{}\mathclose{{}\left|A^{-}}\right|+\mathopen{}\mathclose{{}\left|B^{+}}\right|=\sigma_{a}\mathopen{}\mathclose{{}\left({e}}\right)+n-(\sigma_{b}\mathopen{}\mathclose{{}\left({e}}\right)-1)$. Using that $e$ is $c$-steep, we thus have $\mathopen{}\mathclose{{}\left|A^{-}}\right|+\mathopen{}\mathclose{{}\left|B^{+}}\right|\geq n+cn$. Combining both inequalities, we get $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$.
While steep edges demonstrate how an intuitively common structure cannot be there if assume there are no pivot-edges. However, in randomly generated instances we expect edges not to be steep. Assume we add edges randomly to the graph in the order of their labels. Then, when adding the edge $\mathopen{}\mathclose{{}\left\{v,w}\right\}$, both $v$ and $w$ should already have a similar number of edges. In fact, one can prove that with high probability, there is no $c$-steep in the graph.
Yet, as every $c$-steep edge is pivotable, we want to analyze the graphs without such edges in greater detail. We first look at more simple properties showing that any node only has a limited number of neighbors for which it is at least the $i$-th and at most the $j$-th neighbor.
Lemma 5.13 (Label Spread).
Let $c\in(0,1]$. In a bi-clique $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ without $c$-steep edges, for any $i,j\in[n]$ and $v\in V$ we know that for the label spread $S\coloneqq\mathopen{}\mathclose{{}\left\{v^{\prime}\in N(v)\mid i\leq\sigma_{v^{\prime}}\mathopen{}\mathclose{{}\left({v}}\right)\leq j}\right\}$ it holds that
$\mathopen{}\mathclose{{}\left|S}\right|<j-i+2cn.$
Proof 5.14.
Let $S\coloneqq\mathopen{}\mathclose{{}\left\{v^{\prime}\in N(v)\mid i\leq\sigma_{v^{\prime}}\mathopen{}\mathclose{{}\left({v}}\right)\leq j}\right\}$ and choose a $v^{\prime}\in S$. As the edge $\mathopen{}\mathclose{{}\left\{{v},{v^{\prime}}}\right\}$ is not $c$-steep, we have that $\sigma_{v}\mathopen{}\mathclose{{}\left({v^{\prime}}}\right)\in(i-cn,j+cn)$. Since $\sigma_{v}$ is an injection, we have that $|S|\leq\mathopen{}\mathclose{{}\left|(i-cn,j+cn)}\right|\leq j+cn-(i-cn)=j-i+2cn$.
The next lemma tells us that if a vertex $v$ can reach an edge $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$, $v$ can also reach all edges $e^{\prime}$ incident to either $a$ or $b$ with $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e^{\prime}}}\right)>\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$.
Lemma 5.15.
Given a temporal bi-clique $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$. Fix any $b\in B$ and order all $a_{1},\ldots,a_{n}\in A$ according to $\sigma_{b}$ (i.e., ${a_{i}}\preceq_{b}{a_{i+1}}$ for all $i\in[n-1]$). Then
$\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{1}},{b}}\right\}}}\right)\subseteq\cdots\subseteq\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{n}},{b}}\right\}}}\right)$
and
$\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{1}},{b}}\right\}}}\right)\supseteq\cdots\supseteq\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{n}},{b}}\right\}}}\right).$
The same holds when switching roles for $A$ and $B$.
Proof 5.16.
Consider $v\in\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b}}\right\}}}\right)$ and let $j>i$. Then there exists a temporal path $P$ starting at $v$ and ending in $\mathopen{}\mathclose{{}\left\{{a_{i}},{b}}\right\}$. Assume w.l.o.g. that the path ends in $b$, else remove the latest edge from the path. By adding the edge $\mathopen{}\mathclose{{}\left\{{b},{a_{j}}}\right\}$ to the path, the path stays temporal since ${a_{i}}\preceq_{b}{a_{j}}$ and $v\in\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{j}},{b}}\right\}}}\right)$. The proof for $\operatorname{Out}$ as well as for switched roles of $A$ and $B$ follows analogously.
We can rephrase the last lemma to characterize which vertices remain in $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)$ when replacing $\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$ with $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b}}\right\}$. If $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b}}\right\}$ has a similar label to $\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$, the sets $\operatorname{In}$ and $\operatorname{Out}$ will remain roughly the same.
Corollary 5.17 (Activity Width).
Given a temporal bi-clique $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$. Fix any $b\in B$ and order all $a_{1},\ldots,a_{n}\in A$ according to $\sigma_{b}$ (i.e., ${a_{i}}\preceq_{b}{a_{i+1}}$ for all $i\in[n-1]$).
For every $i\in[n]$ we define $M^{A}_{b,i}\coloneqq\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b}}\right\}}}\right)\cap A$. Then for every $a\in A$ there exists an interval $[k,\ell]$ such that $a\in M^{A}_{b,i}$ if and only if $i\in[k,\ell]$. The same holds for $M^{B}_{b,i}\coloneqq\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b}}\right\}}}\right)\cap B$.
Further assuming that no $c$-pivot-edge exists in $G$, for all $b\in B,i\in[n]$ we have $\mathopen{}\mathclose{{}\left|M^{A}_{b,i}}\right|<cn$, thus on average for each $a\in A$, the interval $[k,\ell]$ such that $a\in M^{A}_{b,i}\iff k\leq i\leq\ell$ has size less than $cn$.
6 Bi-Cliques With Reverted Edges
Now that we have introduced the concept of $c$-pivot-edges and observed some interesting structural properties in bi-cliques which contain no suitable $c$-pivot-edge. The first hope would be that every bi-cliques admits such a $c$-pivot-edge. Unfortunately, this is not the case and we provide a general graph class, the shifted matching graph, in which for every edge $e$, the pivot-set is as small as possible (i.e., $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)=\{a,b\}$) and none of our reduction rules from Section 4 are applicable.
Still, we present a novel technique, called $e$-reverted edges, to construct small spanners for shifted matching graphs which also provides solutions for a significantly larger class of bi-cliques. We relate the technique to $c$-pivot-edges from Section 5 and argue why $e$-reverted edges solve a class distinct from $c$-pivot-edges, beyond the concrete example of shifted matching graphs. Unfortunately, also this is not applicable to all graph classes, which we later demonstrate with a class that we call product graphs in Section 7.
We start by giving the definition of the shifted matching graph.
Definition 6.1 (Shifted Matching Graph).
Let $n\in\operatorname{\mathbb{N}}^{+}$, $A\coloneqq\{a_{0},\ldots,a_{n-1}\}$, and $B\coloneqq\{b_{0},\ldots,b_{n-1}\}$. We define the shifted matching graph on $n$ vertices per side as $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)\coloneqq\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where for every $i,j\in\{0,\ldots,{n}-1\}$ we label the edge $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{a_{i}}},{{b_{j}}}}\right\}}}\right)\coloneqq j-i\operatorname{mod}n.$
Vertex Edge Label
0 1 2 3
$a_{0}$ $b_{0}$ $b_{1}$ $b_{2}$ $b_{3}$
$a_{1}$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{0}$
$a_{2}$ $b_{2}$ $b_{3}$ $b_{0}$ $b_{1}$
$a_{3}$ $b_{3}$ $b_{0}$ $b_{1}$ $b_{2}$
(a) View from $A$
Vertex Edge Label
0 1 2 3
$b_{0}$ $a_{0}$ $a_{3}$ $a_{2}$ $a_{1}$
$b_{1}$ $a_{1}$ $a_{0}$ $a_{3}$ $a_{2}$
$b_{2}$ $a_{2}$ $a_{1}$ $a_{0}$ $a_{3}$
$b_{3}$ $a_{3}$ $a_{2}$ $a_{1}$ $a_{0}$
(b) View from $B$
Figure 4: $\operatorname{SM}\mathopen{}\mathclose{{}\left({4}}\right)$ with vertices $v\in A\sqcup B$ and all their neighbors, ordered by edge label.
Figure 4 shows the shifted matching graph for $n=4$ vertices on both sides. In shifted matching graphs, the earliest and latest neighbors form a matching, thus we cannot apply the reduction from Theorem 4.7. Note that for all $i\in\{0,\ldots,{n}-1\}$, we have $\pi^{-}(a_{i})=b_{i}$ and $\pi^{+}(a_{i})=b_{i-1\operatorname{mod}n}$. Additionally, we observe that each edge at time $t$ has the form $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{i+t\operatorname{mod}n}}}\right\}$. Thus, the edges with label $t$ form a perfect matching and $a_{i}$ sees the $b_{j}$ in ascending order of indices, rotated by $i$ places compared to $a_{0}$.
{observation}
Let $t\in\{0,\dots n-1\}$ be a time-point. Then in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$, the edges with label $t$ form a perfect matching.
We claim that this graph has no $c$-pivot-edges that are suitable for reduction in the sense of Theorem 5.8. To prove that this method fails on the shifted matching graph, we first characterize the set of vertices which can be reached from a vertex $a_{i}$ before and after a given time. Finally, we prove that there, the intersection of $\operatorname{In}$ and $\operatorname{Out}$ for all edges at $a_{i}$ is trivial.
Lemma 6.2.
Let $i,t\in\{0,\ldots,{n}-1\}$ and $a_{i}$ and $b_{i+t\operatorname{mod}n}$ be vertices in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$. Then using edges with label at most $t$, precisely the vertices $a_{i},\dots,a_{i+t\operatorname{mod}n}$ and $b_{i},\dots,b_{i+t\operatorname{mod}n}$ can reach $a_{i}$ and $b_{i+t\operatorname{mod}n}$. Using only edges with label at least $t$, $a_{i}$ and $b_{i+t\operatorname{mod}n}$ can precisely reach the vertices $b_{i+t\operatorname{mod}n},\dots,b_{i-1\operatorname{mod}n}$ and $a_{i+t+1\operatorname{mod}n},\ldots,a_{i}$.
Proof 6.3.
We prove the first part of the statement via induction over $t$. The second part follows analogously by induction over $n-t-1$. For $t=0$, clearly $a_{i}$ as well as $b_{i}$ can only be reached by themselves and each other. By Section 6, we can never compose two edges with label 00.
Now assume the statement holds for all vertices for some $t<n-1$. Thus, precisely $a_{i},\dots,a_{i+t\operatorname{mod}n}$ and $b_{i},\dots,b_{i+t\operatorname{mod}n}$ can reach $a_{i}$ and precisely $a_{i+1\operatorname{mod}n},\dots,a_{i+t+1\operatorname{mod}n}$ and $b_{i+1\operatorname{mod}n},\dots,b_{i+t+1\operatorname{mod}n}$ can reach $b_{i+t+1\operatorname{mod}n}$. Now we look at the situation, when we allow edges with label $t+1$ as well. Since $a_{i}$ and $b_{i+t+1\operatorname{mod}n}$ are connected with an edge with label $t+1$, we can take the union of both sets. No additional vertices can now reach $a_{i}$ and $b_{i+t+1\operatorname{mod}n}$, since after taking an edge with label $t+1$ from a different vertex, we are stuck due to Section 6.
Using Lemma 6.2, we can conclude that using $c$-pivot-edges in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ leads to no reduction.
Lemma 6.4.
For vertices $a_{i},b_{j}$ in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$, we have $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},b_{j}\}$.
Proof 6.5.
Let $t\coloneqq\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{a_{i}}},{{b_{j}}}}\right\}}}\right)=j-i\operatorname{mod}n$ and therefore $b_{j}=b_{i+t\operatorname{mod}n}$. By Lemma 6.2,
$\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},\ldots,a_{i+t\operatorname{mod}n},b_{i\operatorname{mod}n},\ldots,b_{i+t\operatorname{mod}n}\}$
and
$\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i+t+1\operatorname{mod}n},\ldots,a_{i},b_{i+t\operatorname{mod}n},\ldots,b_{i-1\operatorname{mod}n}\}.$
Looking at the intersection, we notice $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},b_{j}\}$.
However, with a new technique we can still construct linear-size spanners for $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$.
For this, we need the concept of $e$-reverted edges. This construction sheds light on different ideas, such as making use of $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ to switch $A$ and $B$, that can be used for other graphs than just $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$. We start by giving the general technique and then proving that it is applicable to construct a linear-size bi-spanner for the shifted matching graph.
Definition 6.6 ($e$-reverted).
[…]
Intuitively, for each edge $e$ there is a linear-size set of edges such that all pairs of vertices $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}$, that share an $e$-reverted edge, are connected via a short path. If ${a^{\prime}}\preceq_{b}{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}$, the path $a^{\prime}\to b\to\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)\to b^{\prime}$ is temporal, since $b^{\prime}$ is also the latest neighbor for $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)$. If ${\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}\preceq_{a}{b^{\prime}}$, the path $a^{\prime}\to\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)\to a\to b^{\prime}$ is temporal, since $a^{\prime}$ is also the earliest neighbor $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)$. See Figure 5.
Figure 5: If at least one of the green or purple paths is temporal, $\{a^{\prime},b^{\prime}\}$ is $\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$-reverted.
Notice that, no matter which $a^{\prime}\in A$ and $b^{\prime}\in B$ we choose, the paths have many edges in common. We always use edges incident to $a$ or $b$ and $\operatorname{\pi^{\text{--}}}$ or $\operatorname{\pi^{+}}$. We use this to construct a small bi-spanner for $D$, if $\operatorname{NotRev}_{{e}}$ is small.
Theorem 6.7.
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be an extremally matched bi-clique of size $n$ and $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}\in{A}\otimes{B}$. Then $D$ has a bi-spanner of size at most $4n-4+\mathopen{}\mathclose{{}\left|\operatorname{NotRev}_{{e}}}\right|$.
Proof 6.8.
Define $S$ to include all edges adjacent to $a$ or $b$ as well as the perfect matchings $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ and all edges in $\operatorname{NotRev}_{{e}}$. First notice that $|S|\leq 4n+\mathopen{}\mathclose{{}\left|\operatorname{NotRev}_{{e}}}\right|$ and we double count at least the first and last edges of $a$ and $b$ (and $e$ if these overlap).
Furthermore, we show that $S$ is a bi-spanner for $D$. Let $a^{\prime}\in A$ and $b^{\prime}\in B$. We distinguish three cases.
Case 1: $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}\in\operatorname{NotRev}_{{e}}$. In this case, the direct edge between $a^{\prime}$ and $b^{\prime}$ is contained in $S$. In all other cases $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}$ is $e$-reverted.
Case 2: ${a^{\prime}}\preceq_{b}{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}$. As argued before, in this case the path $a^{\prime}\to b\to\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)\to b^{\prime}$ is temporal and contained in $S$.
Case 3: ${\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}\preceq_{a}{b^{\prime}}$. In this case, the path $a^{\prime}\to\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)\to a\to b^{\prime}$ is temporal and contained in $S$.
Note that we only require a single edge $e$ to have a small number of non-$e$-reverted edges to obtain a small spanner. Fortunately, the shifted matching graph $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ contains many edges $e$ for which all edges are $e$-reverted.
Lemma 6.9.
Let $n\in\operatorname{\mathbb{N}}^{+}$ and $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$ be an edge with label 00 or $n-1$ in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$. Then $\operatorname{NotRev}_{{e}}=\emptyset$.
Proof 6.10.
We show the statement only for $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)=0$, the proof for label $n-1$ follows analogously. Let wlog. $a=a_{0}$ and $b=b_{0}$ and let $i,j\in\{0,\ldots,{n}-1\}$. To show that $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}$ is $e$-reverted, we distinguish two cases.
Case 1: $i\leq j$. By the structure of the shifted matching graph, we know $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a_{i}}}\right)={b_{i}}\preceq_{a_{0}}{b_{j}}$, and $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}$ is $e$-reverted.
Case 2: $i>j$. In this case, we have ${a_{i}}\preceq_{b_{0}}{a_{j+1\operatorname{mod}n}}=\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b_{j}}}\right)$. Thus, $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}$ is $e$-reverted.
This allows us to apply Theorem 6.7 to $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ to construct a linear-size bi-spanner.
Corollary 6.11.
For every $n\in\operatorname{\mathbb{N}}^{+}$, the graph $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ has a bi-spanner of size $44n-$ 4.
Now we connect the concept of $e$-reverted edges to $c$-pivot-edges from Section 5
Lemma 6.12.
Let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique and $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}\in{A}\otimes{B}$. Then $\mathopen{}\mathclose{{}\left|\operatorname{NotRev}_{{e}}}\right|\geq\mathopen{}\mathclose{{}\left|A\setminus\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\cdot\mathopen{}\mathclose{{}\left|B\setminus\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|$.
Proof 6.13.
Let $b^{\prime}\in B\setminus\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ and $a^{\prime}\in A\setminus\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$. This also implies $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)\notin\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ and $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)\notin\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$. Therefore, we have ${\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}\prec_{b}{a}$ and ${a}\prec_{b}{a^{\prime}}$ and by transitivity of $\prec_{b}$ we get ${\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}\prec_{b}{a^{\prime}}$. Additionally, ${b^{\prime}}\prec_{a}{b}$ and ${b}\prec_{a}{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}$ follows from the definitions of $a^{\prime}$ and $b^{\prime}$. Again, using transitivity we get ${b^{\prime}}\prec_{a}{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}$ and $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}\in\operatorname{NotRev}_{{e}}$.
Lemma 6.12 tells us something about where to look for edges $e$ with many $e$-reverted edges. In fact, only edges that are either among the earliest or the latest edges of the incident vertices are suitable candidates for Theorem 6.7, which are the least likely to be $c$-pivot-edges. This gives us an intuition, why the two techniques can be applied in different situations. This intuition is supported by the concrete example of shifted matching graphs.
7 Product Graphs
In this section we present a general construction to compose two bi-cliques. We show that a careful application yields a bi-clique which cannot be solved by the techniques presented so far. In particular, we present a class of graphs, such that for any edge $e$ the pivot-set of $e$ is sub-linear and the set of non-revertible vertex-pairs grows almost quadratically.
(a) Construction of product graphs. Observe how every edge in $G$ is replaced by a copy of the graph $H$ and every label is the pair of their original labels.
(b) The construction of the path in the proof of Lemma 7.13. Purple labels are the first element of vertex names, orange ones are the second element.
Figure 6: Product graphs and how to find small spanners.
Theorem 7.1.
Let $f\colon\operatorname{\mathbb{N}}\to\operatorname{\mathbb{N}}^{+}$ be any function with $f(n)\in\operatorname{\mathcal{O}}(n)$. There is a set of bi-cliques $\{D_{n}\}_{n\in\operatorname{\mathbb{N}}}=\{(A_{n},B_{n},\lambda_{n})\}_{n\in\operatorname{\mathbb{N}}}$, such that
  1. 1.
    $|A_{n}|=|B_{n}|\in\Theta(n)$;
  2. 2.
    for all $e\in{A_{n}}\otimes{B_{n}}$
    1. (a)
      the size of $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ is in $\operatorname{\mathcal{O}}\mathopen{}\mathclose{{}\left(f(n)}\right)$,
    2. (b)
      the number of the not $e$-reverted edges $\operatorname{NotRev}_{{e}}$ is in $\Omega\mathopen{}\mathclose{{}\left(nf(n)}\right)$.
We do not prove Theorem 7.1 right away. Rather, we first consider an application. For this, choose $f(n)\coloneqq n^{1-\varepsilon}$. Theorem 7.1 tells us that there is a set of graphs $\{D_{n}\}_{n\in\operatorname{\mathbb{N}}}$ such that the size of sides is in $\Theta(n)$, the size of any pivot-set is in $\operatorname{\mathcal{O}}(n^{1-\varepsilon})$, and the number of not reverted edges with respect to any edge is in $\Omega(n^{2-\varepsilon})$.
To prove Theorem 7.1, we define a graph class that we later use to define appropriate $D_{n}$.
Definition 7.2 (Product Graph).
Consider two bi-cliques $G=(A_{G},B_{G},\operatorname{\lambda}_{G})$ and $H=(A_{H},B_{H},\operatorname{\lambda}_{H})$. Define the product graph ${G}\operatorname{\times}{H}=(\mathcal{A},\mathcal{B},\Lambda)$ as
$\mathcal{A}$ $\coloneqq A_{G}\times A_{H},$
$\mathcal{B}$ $\coloneqq B_{G}\times B_{H},$
$\Lambda$ $\coloneqq((a_{G},a_{H}),(b_{G},b_{H}))\mapsto(\operatorname{\lambda}_{G}(a_{G},b_{G}),\operatorname{\lambda}_{H}(a_{H},b_{H})).$
Note that the co-domain of $\Lambda$ is $\operatorname{\mathbb{N}}^{2}$. To make it the proper definition of an edge labeling function, take the lexicographic embedding. For the reader’s benefit, we continue to write tuples. Intuitively, to construct ${G}\operatorname{\times}{H}$, we start with the graph $G$ and replace each vertex of $A_{G}$ with a distinct copy of $A_{H}$ as well as each vertex of $B_{G}$ with a distinct copy of $B_{H}$, see Figure 5(a) for an illustration of this idea. We refer to each such expanded vertex as a bag. Note that this operation is not commutative. The edge labels $\Lambda$ are chosen in such a way that any 2-edge path $x\to y\to z$ with $x$ and $z$ in the same bag is temporal if the respective path in $H$ is temporal and any 2-edge path $x\to y\to z$ with $x$ and $z$ belonging to different bags is temporal if the respective path is temporal in $G$. We now make this rigorous.
Lemma 7.3.
Let $G$ and $H$ be bi-cliques. Let $P=(g_{1},h_{1})(g_{2},h_{2})\dots(g_{\ell},h_{\ell})$ be a temporal path in ${G}\operatorname{\times}{H}$. Then
  1. 1.
    the sequence $g_{1}g_{2}\dots g_{\ell}$ is a temporal path in $G$;
  2. 2.
    if $\operatorname{\lambda}_{G}(g_{1},g_{2})=\operatorname{\lambda}_{G}(g_{\ell-1},g_{\ell})$, the sequence $h_{1}h_{2}\dots h_{\ell}$ is a temporal path in $H$.
Proof 7.4.
By choice of $\Lambda$, for all $i\in[\ell-2]$ we have $\operatorname{\lambda}_{G}(g_{i},g_{i+1})\leq\operatorname{\lambda}_{G}(g_{i+1},g_{i+2})$ and so $g_{1}g_{2}\dots g_{\ell}$ is temporal in $G$. Additionally, if $\operatorname{\lambda}_{G}(g_{1},g_{2})=\operatorname{\lambda}_{G}(g_{\ell-1},g_{\ell})$, for all $i\in[\ell-1]$ the values $\operatorname{\lambda}_{G}(g_{i},g_{i+1})$ are equal. By lexicographic ordering, for all $i\in[\ell-2]$ we have $\operatorname{\lambda}_{H}(h_{i},h_{i+1})\leq\operatorname{\lambda}_{H}(h_{i+1},h_{i+2})$ and so $h_{1}h_{2}\dots h_{\ell}$ is temporal in $H$.
This gives us enough tools to prove how the pivot-set of an edge $e=\mathopen{}\mathclose{{}\left\{{(u_{G},u_{H})},{(v_{G},v_{H})}}\right\}\in E({G}\operatorname{\times}{H})$ in ${G}\operatorname{\times}{H}$ relates to the pivot-set of $\mathopen{}\mathclose{{}\left\{{u_{G}},{v_{G}}}\right\}$ in $G$. For this, define $\operatorname{\phi}\colon V({G}\operatorname{\times}{H})\to V(G)$ as $(u_{G},u_{H})\overset{\operatorname{\phi}}{\mapsto}u_{G}$. This denotes the function that yields for every vertex in ${G}\operatorname{\times}{H}$ its bag, namely, its original vertex in $G$. We will also naturally extend $\operatorname{\phi}$ to edges (i.e., $\operatorname{\phi}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{(u_{G},u_{H})},{(v_{G},v_{H})}}\right\}}}\right)=\mathopen{}\mathclose{{}\left\{{u_{G}},{v_{G}}}\right\}$). Intuitively, for any vertex $v$ in the pivot-set of an edge $e\in E({G}\operatorname{\times}{H})$, the vertex $\operatorname{\phi}\mathopen{}\mathclose{{}\left({v}}\right)$ must be in the pivot-set of $\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)$ in $G$.
Lemma 7.5.
Let $G$ and $H$ be bi-cliques and let $e\in E({G}\operatorname{\times}{H})$. We have
  1. 1.
    $\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\subseteq\operatorname{In}\mathopen{}\mathclose{{}\left({\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)$,
  2. 2.
    $\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\subseteq\operatorname{Out}\mathopen{}\mathclose{{}\left({\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)$, and
  3. 3.
    $\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\subseteq\operatorname{In}\mathopen{}\mathclose{{}\left({\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)$.
Proof 7.6.
The third statement follows from the first two by
$\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\subseteq\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\cap\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\subseteq\operatorname{In}\mathopen{}\mathclose{{}\left({\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right)}}\right).$
We prove the first statement, the second is analogous. Consider any $e=\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}\in E({G}\operatorname{\times}{H})$ and $w\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$. We need to show that there is a temporal path from $\operatorname{\phi}\mathopen{}\mathclose{{}\left({w}}\right)$ to one of $\operatorname{\phi}\mathopen{}\mathclose{{}\left({u}}\right)$ or $\operatorname{\phi}\mathopen{}\mathclose{{}\left({v}}\right)$ arriving not later than $\operatorname{\lambda}_{G}(\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right))$. For this consider the path $P=p_{1}p_{2}\dots p_{\ell}$ from $w$ to $e$ that does not end later than $\Lambda(e)$. By Lemma 7.3, the path $P^{\prime}\coloneqq\operatorname{\phi}\mathopen{}\mathclose{{}\left({p_{1}}}\right)\operatorname{\phi}\mathopen{}\mathclose{{}\left({p_{2}}}\right)\dots\operatorname{\phi}\mathopen{}\mathclose{{}\left({p_{\ell}}}\right)$ is temporal in $G$. Note that, $\operatorname{\phi}\mathopen{}\mathclose{{}\left({p_{1}}}\right)=\operatorname{\phi}\mathopen{}\mathclose{{}\left({w}}\right)$ and $\operatorname{\phi}\mathopen{}\mathclose{{}\left({p_{\ell}}}\right)\in\mathopen{}\mathclose{{}\left\{\operatorname{\phi}\mathopen{}\mathclose{{}\left({u}}\right),\operatorname{\phi}\mathopen{}\mathclose{{}\left({v}}\right)}\right\}$. For the last edge $e^{\prime}=p_{\ell-1}p_{\ell}$ of $P$, we have $\Lambda(e^{\prime})\leq\Lambda(e)$, and so $\operatorname{\lambda}_{G}(\operatorname{\phi}\mathopen{}\mathclose{{}\left({e^{\prime}}}\right))\leq\operatorname{\lambda}_{G}(\operatorname{\phi}\mathopen{}\mathclose{{}\left({e}}\right))$. Thus, $P^{\prime}$ is a path we were searching for.
We now focus on product graphs, where both $G$ and $H$ are shifted matching graphs, as we will use these to prove Theorem 7.1. Therefore, they will serve as the family of graphs which the already introduced techniques do not solve.
Definition 7.7.
Let $m,k\in\operatorname{\mathbb{N}}$, Define the graph $\operatorname{SM}(m,k)={\operatorname{SM}\mathopen{}\mathclose{{}\left({m}}\right)}\operatorname{\times}{\operatorname{SM}\mathopen{}\mathclose{{}\left({k}}\right)}$ to be the product graph obtained by taking $\operatorname{SM}\mathopen{}\mathclose{{}\left({m}}\right)$ as the outer graph and $\operatorname{SM}\mathopen{}\mathclose{{}\left({k}}\right)$ as the inner graph.
Observe that each side of $\operatorname{SM}(m,k)$ has $m$ bags each of size $k$ and so, overall, $\operatorname{SM}(m,k)$ has $mk$ vertices per side. We now investigate the size of the pivot-sets in $\operatorname{SM}(m,k)$ as well as the number of the non-reverted edges.
Lemma 7.8.
Let $m,k\in\operatorname{\mathbb{N}}$ and consider $\operatorname{SM}(m,k)$. For any $e\in E(\operatorname{SM}(m,k))$, we have that the size of the pivot-set $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ is at most $2k$.
Proof 7.9.
Let $e=\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}\in E(\operatorname{SM}(m,k))$. By Lemmas 6.4 and 7.5, we have $\operatorname{\phi}\mathopen{}\mathclose{{}\left({\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}}\right)\subseteq\mathopen{}\mathclose{{}\left\{\operatorname{\phi}\mathopen{}\mathclose{{}\left({u}}\right),\operatorname{\phi}\mathopen{}\mathclose{{}\left({v}}\right)}\right\}$ and therefore every $w\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ must be in the same bag as $u$ or $v$. As the size of each bag is $k$, we have $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\leq 2k$.
Lemma 7.10.
Let $k,m\in\operatorname{\mathbb{N}}$ and consider $\operatorname{SM}(m,k)$. For any $e\in E(\operatorname{SM}(m,k))$, the size of the set of not $e$-reverted edges $\operatorname{NotRev}_{{e}}$ is at least $(m-1)\binom{k}{2}$.
Proof 7.11.
Recall, that there is an order on the super vertices $a_{0},a_{1},\dots,a_{m-1}$ and $b_{0},b_{1},\dots,b_{m-1}$ such that for all $a_{i}\in A_{G}$ and $b_{j}\in B_{G}$ we have $\operatorname{\lambda}_{G}(\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\})=(j-i)\operatorname{mod}m$. For all $0\leq i<m$, we write $A_{i}\coloneqq\operatorname{\phi}^{-1}(a_{i})$ and $B_{j}\coloneqq\operatorname{\phi}^{-1}(b_{i})$.
Now consider any edge $e\in E(\operatorname{SM}(m,k))$. Notice that for all $t\in\operatorname{\mathbb{Z}}$, the temporal graph $\operatorname{SM}(m,k)$ is isomorphic to itself (in the sense that the map does not change the labels between two vertices) with the function
$(a_{i},a_{j})\mapsto(a_{(i+t)\operatorname{mod}m},a_{j})\text{ and }(b_{i},b_{j})\mapsto(b_{(i+t)\operatorname{mod}m},b_{j}).$
We can therefore assume without loss of generality, that $e=\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$ with $a\in A_{0}$ and let $j$ be such that $b\in B_{j}$. Therefore, for any $0\leq x<y<m$, $b_{x}\in B_{x}$, and $b_{y}\in B_{y}$, we have ${b_{x}}\prec_{a}{b_{y}}$.
One can verify, that $\operatorname{SM}(m,k)$ is extremally matched. So, we obtain that
[…]
Now consider any $11\leq\ell\leq m-$ 1 as well as $a^{\prime},a^{\prime\prime}\in A_{\ell}$ with ${a^{\prime}}\succ_{b}{a^{\prime\prime}}$. By the properties of shifted matching graphs, $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)\in B_{\ell}$ and $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({a^{\prime\prime}}}\right)\in B_{(\ell-1)\operatorname{mod}m}=B_{\ell-1}$. By choice of $a$, we have ${\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({a^{\prime}}}\right)}\succ_{a}{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({a^{\prime\prime}}}\right)}$. So, $(a^{\prime},\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({a^{\prime\prime}}}\right))\in\operatorname{NotRev}_{{e}}$.
For every $11\leq\ell\leq m-$ 1, there are $\binom{k}{2}$ distinct pairs contributing to $\operatorname{NotRev}_{{e}}$. Additionally, for every of the $m-1$ choice of $\ell$, these pairs are different. Overall, $\mathopen{}\mathclose{{}\left|\operatorname{NotRev}_{{e}}}\right|\geq(m-1)\binom{k}{2}$.
Now we can finally conclude that these two lemmas together mean that our tools are not strong enough to solve this graph class yet.
Proof 7.12 (Proof of Theorem 7.1).
Let $m_{n}\coloneqq 1+\mathopen{}\mathclose{{}\left\lceil\frac{n}{f(n)}}\right\rceil$ and $k_{n}\coloneqq 1+f(n)$ and choose $G_{n}=\operatorname{SM}(m_{n},k_{n})$. Denote with $s_{n}=m_{n}k_{n}$ the size of each side of $G_{n}$. We have
$s_{n}$ $=m_{n}k_{n}=\mathopen{}\mathclose{{}\left(1+\mathopen{}\mathclose{{}\left\lceil\frac{n}{f(n)}}\right\rceil}\right)\mathopen{}\mathclose{{}\left(1+f(n)}\right)\geq\frac{n}{f(n)}\cdot f(n)\geq n,$
and
$s_{n}$ $\leq\mathopen{}\mathclose{{}\left(2+\frac{n}{f(n)}}\right)\cdot(1+f(n))\leq 2+2f(n)+2n.$
Since $f(n)\in\operatorname{\mathcal{O}}(n)$, we have $s_{n}\in\Theta(n)$. By Lemma 7.8, we have that any pivot set is in $\operatorname{\mathcal{O}}(k_{n})=\operatorname{\mathcal{O}}(f(n))$, and by Lemma 7.10 any set of not reverted edges has size at least
$(m_{n}-1)\binom{k_{n}}{2}$ $\geq\frac{n}{f(n)}\mathopen{}\mathclose{{}\left(\frac{(f(n)+1)\cdot f(n)}{2}}\right)\geq\frac{nf(n)}{2}.$
With our previous techniques, we are not able to give a linear-size spanner for every product graph. By Theorem 7.1, for $f(n)\in\Theta(\sqrt{n})$, Theorem 5.6 and Theorem 6.7 only give us $\Theta(n^{3/2})$ spanners. Still, product graphs exhibit a lot of structure, in the sense that all edges between two bags have roughly the same time labels. We use this to construct bi-spanners for the product graph based on bi-spanners for the two underlying graphs.
Lemma 7.13.
Let $G=\mathopen{}\mathclose{{}\left({A_{G}},{B_{G}},{\operatorname{\lambda}_{G}}}\right)$ and $H=\mathopen{}\mathclose{{}\left({A_{H}},{B_{H}},{\operatorname{\lambda}_{H}}}\right)$ be temporal bi-cliques of sizes $n_{G}$ and $n_{H}$ and let $S_{G},S_{H}$ be temporal bi-spanners respectively. Then, there is a bi-spanner in ${G}\operatorname{\times}{H}$ of size at most $|S_{G}|n_{H}+|S_{H}|n_{G}$.
Proof 7.14.
Let $b_{G}\in B_{G}$. Define $\pi^{+}_{{S_{G}}}({b_{G}})$ to be the $a_{G}\in A_{G}$, such that $\mathopen{}\mathclose{{}\left\{{a_{G}},{b_{G}}}\right\}$ is the latest edge in $S_{G}$ that is incident to $b_{G}$. Note that since $S_{G}$ is a spanner, every $b_{G}$ must be incident to at least one edge. Furthermore, choose $b_{H}^{*}\in B_{H}$ and $a_{H}^{*}\in A_{H}$ such that the label of $\mathopen{}\mathclose{{}\left\{{a_{H}^{*}},{b_{H}^{*}}}\right\}$ is minimal.
We construct a temporal spanner $S$ for ${G}\operatorname{\times}{H}$ as follows. For every $\mathopen{}\mathclose{{}\left\{{a_{G}},{b_{G}}}\right\}\in S_{G}$ and $a_{H}\in A_{H}$, add $\mathopen{}\mathclose{{}\left\{{(a_{G},a_{H})},{(b_{G},b_{H}^{*})}}\right\}$ to $S$. Additionally, for every $\mathopen{}\mathclose{{}\left\{{a_{H}},{b_{H}}}\right\}\in S_{H}$ and $b_{G}\in B_{G}$, add $\mathopen{}\mathclose{{}\left\{{(\pi^{+}_{{S_{G}}}({b_{G}}),a_{H})},{(b_{G},b_{H})}}\right\}$ to $S$.
We can immediately verify that this results in a set of size $|S|\leq|S_{G}|n_{H}+|S_{H}|n_{G}$. To show temporal connectivity, let $(a_{G},a_{H})\in A_{G}\times A_{H}$ and $(b_{G},b_{H})\in B_{G}\times B_{H}$, we aim to construct a path $P$ from $(a_{G},a_{H})$ to $(b_{G},b_{H})$. Since $S_{G}$ is a bi-spanner for $G$, there is a temporal path $P_{G}$ from $a_{G}$ to $b_{G}$ in $S_{G}$. For every edge $\mathopen{}\mathclose{{}\left\{{a_{G}^{P}},{b_{G}^{P}}}\right\}$ in $P_{G}$ except for the last one, we append $\mathopen{}\mathclose{{}\left\{{(a_{G}^{P},a_{H})},{(b_{G}^{P},b_{H}^{*})}}\right\}$ to $P$. Note that these edges are in $S$, since $\mathopen{}\mathclose{{}\left\{{a_{G}^{P}},{b_{G}^{P}}}\right\}\in S_{G}$. Furthermore, it is easy to verify that two neighboring edges share an endpoint and since $P_{G}$ is temporal, path $P$ is temporal in ${G}\operatorname{\times}{H}$. Since we left out exactly the last edge of $P_{G}$, the path $P$ ends in some vertex $(a_{G}^{\prime},a_{H})$. We distinguish two cases.
Case 1: $a_{G}^{\prime}=\pi^{+}_{{S_{G}}}({b_{G}})$. As $S_{H}$ is a bi-spanner for $H$, there exists a temporal path $P_{H}$ from $a_{H}$ to $b_{H}$ in $S_{H}$. For an edge $\mathopen{}\mathclose{{}\left\{{a_{H}^{P}},{b_{H}^{P}}}\right\}$ in $P_{H}$, append $\mathopen{}\mathclose{{}\left\{{(a_{G}^{\prime},a_{H}^{P})},{(b_{G},b_{H}^{P})}}\right\}$ to $P$. Since $a_{G}^{\prime}=\pi^{+}_{{S_{G}}}({b_{G}})$, this edge is in $S$ by construction. Furthermore, these edges have a higher label than all previous edges in $P$, since $\mathopen{}\mathclose{{}\left\{{a_{G}^{\prime}},{b_{G}}}\right\}$ must be the last edge of $P_{G}$. As $P_{H}$ is temporal, path $P$ remains temporal and it is easy to verify that $P$ is indeed a path from $(a_{G},a_{H})$ to $(b_{G},b_{H})$.
[…]
If $G$ and $H$ each have linear-size bi-spanners, this proves the existence of a linear-size bi-spanner for ${G}\operatorname{\times}{H}$. Since we know from Section 6 how to construct a linear-size bi-spanner for $\operatorname{SM}\mathopen{}\mathclose{{}\left({m}}\right)$, we can now construct a linear-size bi-spanner for $\operatorname{SM}(m,k)$.
Additionally, we can augment Lemma 7.13 to also deal with different graphs $H_{e}$ per edge $e$ in $G$. To do this, we have to change the definition of our representatives from $a_{H}^{*}$ and $b_{H}^{*}$ to functions $r_{A},r_{B}$, such that for $b_{G}\in B_{G}$ the label of $\mathopen{}\mathclose{{}\left\{{(\pi^{+}_{{S_{G}}}({b_{G}}),r_{A}(b_{G}))},{(b_{G},r_{B}(b_{G}))}}\right\}$ is minimal. This enables us not only to construct bi-spanners for $G\times H$ but any bi-cliques that can be decomposed into bags of vertices.
8 Conclusion
Figure 7: Overview of techniques. Previously proposed techniques are blue, new ones are green. Example families of graphs are purple. The red area contains instances that are not solved by any known techniques.
We have made a significant step forwards towards answering whether temporal cliques admit linear-size temporal spanners, see Figure 7 for an illustration. The obvious problem that stems from this paper, is whether classes of graphs exist that are not solved by our techniques. We conjecture that this is the case but the classes of graphs remaining should be quite artificial. Another possible avenue is to generalize some of our techniques to include a wider variety of graphs. For example, we could consider investigating the temporal ordering of edges and try to generalize the concept of reverted edges. Also, product graphs and their solution suggest a way in which temporal bi-cliques may be composed and decomposed. Future research could generalize this idea and more closely analyze the cases in which a temporal spanner can be obtained by combining temporal spanners on subinstances. Towards the opposite direction, it is now easier to find a $\Omega(n\log n)$ counterexample, if it exists at all, since we have provided many properties that cannot exist in a counterexample.

### v2

@@
-Towards Linear Spanners in All Temporal Cliques
+How to Reduce Temporal Cliques to Find Sparse Spanners
 Abstract
-Many real-world networks, like transportation networks and social networks, are dynamic in the sense that the edge set may change over time, but these changes are known in advance.
+Many real-world networks, such as transportation or trade networks, are dynamic in the sense that the edge-set may change over time, but these changes are known in advance.
 This behavior is captured by the temporal graphs model, which has recently become a trending topic in theoretical computer science.
@@
 So far, the best known result is the existence of temporal spanners with $\operatorname{\mathcal{O}}(n\log n)$ many edges.
-We present significant progress towards proving that linear-size temporal spanners exist in all temporal cliques.
-We adapt techniques used in previous works and heavily expand and generalize them to provide a simpler and more intuitive proof of the $\operatorname{\mathcal{O}}(n\log n)$ bound.
-Moreover, we use our novel approach to show that a large class of temporal cliques, called edge-pivot graphs, admit linear-size temporal spanners.
-To contrast this, we investigate other classes of temporal cliques that do not belong to the class of edge-pivot graphs.
-We introduce two such graph classes and we develop novel techniques for establishing the existence of linear temporal spanners in these graph classes as well.
+We present significant progress towards proving whether linear-size temporal spanners exist in all temporal cliques.
+We adapt techniques used in previous works and heavily expand and generalize them.
+This allows us to show that the existence of a linear spanner in cliques and bi-cliques is equivalent and using this, we provide a simpler and more intuitive proof of the $\operatorname{\mathcal{O}}(n\log n)$ bound by giving an efficient algorithm for finding linearithmic spanners.
+Moreover, we use our novel and efficiently computable approach to show that a large class of temporal cliques, called edge-pivotable graphs, admit linear-size temporal spanners.
+To contrast this, we investigate other classes of temporal cliques that do not belong to the class of edge-pivotable graphs.
+We introduce two such graph classes and we develop novel algorithmic techniques for establishing the existence of linear temporal spanners in these graph classes as well.
 keywords:
@@
 Many real-world networks, like transportation networks with scheduled train connections or social networks with repeating meeting schedules, are dynamic, i.e., their edge-set can change over time.
-To address this, recent research in theoretical computer science has shifted from static graphs to temporal graphs [15, 18].
+To address this, temporal graphs have gained significant attention in recent research in theoretical computer science  [15, 18].
 In these graphs it is assumed that the edge-set can change in every time step but the schedule of the availability of edges is known in advance.
@@
 A general theme is that insights from the static graph setting are only of limited value in temporal graphs, since the temporal availability of the edges makes these problems much harder or even infeasible to solve.
-For example, even fundamental theorems like Menger’s Theorem do not hold for temporal graphs [17].
-Also, one of the most fundamental problems in graphs, that of finding a sparse spanner, is much harder in temporal graphs.
+For example, even fundamental statements like Menger’s Theorem do not hold for temporal graphs [17].
+Also, one of the most essential problems in graphs, that of finding a sparse spanner, is much harder in temporal graphs.
 A sparse spanner is a small subset of the edge-set that ensures all-pairs connectivity.
 In static graphs, it is well-known that linear-size spanners exist and can be computed efficiently, e.g., minimum spanning trees [8].
-Finding sparse temporal spanners is much more intricate in temporal graphs, since there paths do not necessarily compose.
+Finding sparse temporal spanners is much more intricate in temporal graphs, since there paths do not necessarily compose (i.e., concatenating two paths where the endpoint of the first is the starting point of the second does not necessarily yield a temporal path).
 More specifically, Kempe, Kleinberg, and Kumar [15] showed that there is a class of temporal graphs having $\Omega(n\log n)$ many edges, such that no single edge can be removed while preserving temporal connectivity.
@@
 Not much progress has been made since then, with most work on the topic veering off towards the study of temporal spanners in random temporal graphs [7], or focusing on low stretch [3, 4] or a game-theoretic setting [5].
-In this paper, we present significant progress on the long-standing open problem by showing that a large class of temporal cliques admits linear-size temporal spanners.
+In this paper, we present significant progress on the long-standing open problem by showing that a large class of temporal cliques does admit linear-size temporal spanners.
 1.1 Our Contribution
@@
 In particular, in [6] a temporal spanner in cliques is computed by a reduction from a clique to a bi-clique.
-We deepen this connection by showing that the existence of a linear spanner in cliques and bi-cliques is equivalent.
+We deepen this connection by showing that the existence of a linear spanner in cliques and bi-cliques actually is equivalent.
 Furthermore, in the reduction of Casteigts, Peters, and Schoeters [6], the authors exploit the property that the set containing the earliest edge of each vertex is a perfect matching of the bi-clique and that this also holds for the set of latest edges.
 We show that this property can also be assumed when searching for bi-spanners in bi-cliques directly.
-This yields a simpler and more intuitive proof, that temporal cliques admit temporal spanners of size $O(n\log n)$.
+This yields a simple and more intuitive algorithm for computing linearithmic spanners, thus proving the existence of size $O(n\log n)$ temporal spanners in temporal cliques.
 Moreover, we reconsider the notion of a pivot-vertex [6, 7].
-This is a vertex $u$ that can be reached by every vertex in the graph until some time step $t$, and then $u$ can reach every other vertex in the graph.
-We transfer this notion to bi-cliques, extend it to pivot-edges, and we strengthen it significantly to find small spanners.
-If a bi-clique contains a pivot-edge, we can iteratively reduce the size of the graph, while maintaining a linear-size temporal spanner, until the instance is solved, or the instance no longer has a pivot-edge.
-This shows that a large class of graphs admits linear-size temporal spanners.
-In the second part of the paper, we investigate temporal cliques that are not edge-pivot graphs.
-We found two such graph classes, called shifted matching graphs and product graphs, and for both we provide novel techniques that allow us to show that linear-size spanners always exist.
-Thus we extend the class of temporal cliques that admit linear-size temporal spanners even further.
+This is a vertex $u$ that can be reached by every vertex in the graph until some time step $t$, and then $u$ can reach every other vertex in the graph by a temporal path starting at or after time $t$.
+We transfer this notion to bi-cliques, extend it to $c$-pivot-edges, and we strengthen it significantly to find small spanners.
+If a bi-clique contains a $c$-pivot-edge, we can iteratively reduce the size of the graph, while maintaining a linear-size temporal spanner, until the instance is solved, or the instance no longer has a $c$-pivot-edge.
+This reduction rule is widely applicable and yields linear-size temporal spanners in many cases.
+In the second part of the paper, we investigate temporal cliques that are not edge-pivotable graphs.
+We identify two such graph classes, called shifted matching graphs and product graphs, and for both we provide novel techniques that allow us to efficiently compute linear-size spanners.
+Thus, we extend the class of temporal cliques that admit linear-size temporal spanners even further.
 It is open if our classes cover every temporal clique.
@@
 The study of temporal spanners was initiated by Kempe, Kleinberg and Kumar [15].
-In their seminal paper, they introduced the temporal graphs model and showed that spanners on temporal hypercubes have $\Omega(n\log n)$ edges.
-Axiotis and Fotakis [2] improved on this by providing a class of graphs with $\Theta(n^{2})$ edges, such that no edge can be removed while maintaining temporal connectivity.
-Later, Casteigts, Peters, and Schoeters [6] studied temporal cliques and showed that temporal spanners of size $O(n\log n)$ always exist.
+In their seminal paper, they introduce the temporal graphs model and show that spanners on temporal hypercubes have $\Omega(n\log n)$ edges.
+Axiotis and Fotakis [2] improve on this by providing a class of graphs with $\Theta(n^{2})$ edges, such that no edge can be removed while maintaining temporal connectivity.
+Later, Casteigts, Peters, and Schoeters [6] study temporal cliques and show that temporal spanners of size $O(n\log n)$ always exist.
 Their "fireworks" algorithm combines different techniques, such as dismountability and delegation, to reduce the clique to a bipartite graph, and then compute a spanner on that graph, that can then be translated to a spanner of the original graph.
-Casteigts, Raskin, Renken and Zamaraev [7] took a different approach.
-They considered random temporal graphs, where the host graph is a clique, and each edge is added to the temporal graph with probability $p$.
-They showed a sharp threshold for the generated temporal graph being temporally connected, and, even more surprisingly, they showed that the generated graph will contain an almost optimal spanner of size $2n+o(n)$.
-On a different note, Bilò, D’Angelo, Gualà, Leucci, and Rossi [4] studied whether temporal cliques admit small temporal spanners with low stretch.
+Casteigts, Raskin, Renken and Zamaraev [7] take a different approach.
+They consider random temporal graphs, where the host graph is a clique, and each edge is added to the temporal graph with probability $p$.
+They show a sharp threshold for the generated temporal graph being temporally connected.
+Even more surprisingly, they show that as soon as this threshold is reached, the generated graph will contain an almost optimal spanner of size $2n+o(n)$.
+On a different note, Bilò, D’Angelo, Gualà, Leucci, and Rossi [4] investigate whether temporal cliques admit small temporal spanners with low stretch.
 Their main result shows that they can always compute a temporal spanner with stretch $O(\log n)$ and $O(n\log^{2}n)$ edges.
@@
 For $n\in\operatorname{\mathbb{N}}^{+}$, let $[n]=\{1,\dots,n\}$.
-Let $(V,E)$ be an undirected graph.
-This together with a labeling function $\operatorname{\lambda}\colon E\to\operatorname{\mathbb{N}}$ forms the temporal graph $G=(V,E,\operatorname{\lambda})$.
-Semantically, an edge $e\in E$ is present at time $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$.
-For a vertex-set $U\subseteq V$, write $G[U]$ for the subgraph of $G$ induced by $U$.
-For $v\in V$, let $N(v)$ denote the set of neighbors of $v$, that is, all vertices $u$ with $\mathopen{}\mathclose{{}\left\{{v},{u}}\right\}\in E$.
-For $v\in V$ and $u,u^{\prime}\in N(v)$, we define the order ${u}\preceq_{v}{u^{\prime}}$, if the edge $\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}$ is not later than the edge $\mathopen{}\mathclose{{}\left\{{u^{\prime}},{v}}\right\}$, i.e., if $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u}},{{v}}}\right\}}}\right)\leq\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u^{\prime}}},{{v}}}\right\}}}\right)$.
-We define $\prec_{v}$, $\succeq_{v}$, and $\succ_{v}$ analogously.
+[…]
 […]
@@
 If both sets $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ form a perfect matching for $V$, then $G$ is extremally matched.
-Also, for $v,w\in V$, and $e=\mathopen{}\mathclose{{}\left\{{v},{w}}\right\}\in E$, we write $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)\in[{n}]$ for the index of $e$, when ordering all edges incident to $v$ ascendingly by their label (e.g., if $w=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)$, then $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)=1$).
+Also, for $v,w\in V$, and $e=\mathopen{}\mathclose{{}\left\{{v},{w}}\right\}\in E$, we write $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)$ or simply $\sigma_{v}\mathopen{}\mathclose{{}\left({w}}\right)$ for the index of $e$ in $[{n}]$, when ordering all edges incident to $v$ ascendingly by their label (e.g., if $w=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)$, then $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)=1$).
 Let $v\in V$ be a vertex and $t\in\operatorname{\mathbb{N}}$ be a time step.
@@
 In these cases, the set of edges is self-evident and for cliques we simply omit it and write $C=(V,\operatorname{\lambda})$.
-We, similarly, denote complete bipartite temporal graphs by a triple $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $A$ and $B$ are the two partitions and $\operatorname{\lambda}\colon{A}\otimes{B}\to\operatorname{\mathbb{N}}$ is the corresponding labeling function.
+We, similarly, denote complete bipartite temporal graphs by a triple $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $A$ and $B$ are the two parts and $\operatorname{\lambda}\colon{A}\otimes{B}\to\operatorname{\mathbb{N}}$ is the corresponding labeling function.
 When necessary, we refer to the set of edges as $E(C)\coloneqq{V}\otimes{V}$ or $E(D)\coloneqq{A}\otimes{B}$.
 We call $C$ a temporal clique and $D$ a temporal bi-clique.
-Finally, we define a directed analogue of spanners for temporal bi-cliques.
+Finally, we define an asymmetric analogue of spanners for temporal bi-cliques.
 We call a set $S\subseteq{A}\otimes{B}$ a bi-spanner for $D$, if every $a\in A$ can reach every $b\in B$ through a temporal path that is contained in $S$.
+Note that here, only one side needs to reach the other.
 For $e=\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}$ in temporal cliques or bi-cliques, notice that all vertices must be in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ or $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$, since all vertices have an edge to $u$ or $v$, which puts this vertex in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ or $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$, depending on the time-label.
@@
 (a) Blue edges and paths have labels earlier than $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$, red ones are later.
-(b) A bi-clique where $a_{1}$ cannot reach $a_{2}$ after time $2$ or can be reached by $a_{2}$ until time $4$, so $\operatorname{In}_{v}^{3}\cup\operatorname{Out}_{v}^{3}\neq V$.
+(b) A bi-clique where $a_{1}$ cannot reach $a_{2}$ using a path starting after time $2$ and $a_{1}$ can not be reached by $a_{2}$ using a path arriving before $5$, so $\operatorname{In}_{a_{1}}^{3}\cup\operatorname{Out}_{a_{1}}^{3}\neq V$.
 Figure 1: Examples for the sets $\operatorname{In}$ and $\operatorname{Out}$.
@@
 Note that for the case of bi-cliques considering both endpoints of $e$ is crucial to ensure that every vertex has a direct edge to an endpoint.
-Otherwise, we simple examples exist having no vertex $v$ with time $t$ with $\operatorname{In}_{v}^{t}\cup\operatorname{Out}_{v}^{t}=A\sqcup B$.
+Otherwise, there exist simple instances with a vertex $v$ and a time $t$ with $\operatorname{In}_{v}^{t}\cup\operatorname{Out}_{v}^{t}\neq A\sqcup B$.
 See Figure 1(b).
@@
 Analogously, let $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that any bi-clique in which both sides have $n$ vertices admits a bi-spanner with at most $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges.
-We study the open problem whether $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ holds.
+We study the open problem whether $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ and whether $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ holds.
 3 Cliques and Bi-Cliques
@@
 The worst-case minimum size of a spanner for temporal cliques and temporal bi-cliques only differs by a constant factor, that is, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\Theta(\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right))$.
-We start with a technical lemma that allows us to focus on injective labeling functions for the rest of the paper. Note that the statement holds for all temporal graphs.
+We start with a technical lemma that allows us to focus on globally injective labeling functions for the rest of the paper.
+This slightly extends a result by [6] which proved the generality of locally injective labelings.
+Note that the statement holds for all temporal graphs.
 Lemma 3.2.
@@
 Then there is an injective labeling function $\lambda^{\prime}$ s.t.
-every spanner $S\subseteq E$ for $(V,E,\lambda^{\prime})$ is also a a spanner for the original graph $G$.
+every spanner $S\subseteq E$ for $(V,E,\lambda^{\prime})$ is also a spanner for the original graph $G$.
 In case the original graph $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ is a bi-clique, the same holds for a bi-spanner $S$.
@@
 We now show that $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
-To this end, we take a bi-clique and add edges within the partitions that yield no benefit in the construction of a spanner.
+To this end, we take a bi-clique and add edges within the parts that yield no benefit in the construction of a spanner.
 Lemma 3.8.
@@
 As $S$ is a spanner, there is a temporal $ab$-path $P\subseteq S$.
-If $P$ contains any edges within $A$, no edge to $B$ can be used afterward, due to the labeling.
+If $P$ contains any edges within $A$, no edge to $B$ can be used afterwards.
 Similarly, once the path arrives at $B$, it is already too late to use any edge within $B$.
@@
 By Lemma 3.8, there is a temporal clique $C=(A\sqcup B,\operatorname{\lambda}^{\prime})$ such that any spanner $S$ in $C$ satisfies $\mathopen{}\mathclose{{}\left|S}\right|\geq\mathopen{}\mathclose{{}\left|S^{*}}\right|$.
-Since the vertex set of $C$ has size $2n$, there is a spanner of size $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)$ in $C$, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
+Since the vertex set of $C$ has size $2n$, there is a spanner of size at most $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)$ in $C$, and thus we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
 To conclude this section, we just need to show, that $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\geq\frac{1}{4}\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({2n}}\right)$.
@@
 Let $n\in\operatorname{\mathbb{N}}$ and let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique of size $n$.
-To construct a bi-clique of size $n+1$, we clone one vertex on each side.
+We construct a bi-clique of size $n+1$ with at least the same minimum spanner size as $D$ by cloning one vertex on each side.
 To be exact, take $a\in A$ and $b\in B$ and let
@@
 $B^{\prime}$ $\coloneqq B\cup\{b^{\prime}\},$
-[…]
+$\operatorname{\lambda}^{\prime}$ $\coloneqq x,y\mapsto\begin{cases}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{a}},{{b}}}\right\}}}\right),\quad&\text{if }x=a^{\prime},y=b^{\prime},\\ \operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{a}},{{y}}}\right\}}}\right),\quad&\text{else if }x=a^{\prime},\\ \operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{x}},{{b}}}\right\}}}\right),\quad&\text{else if }y=b^{\prime},\\ \operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{x}},{{y}}}\right\}}}\right),\quad&\text{otherwise},\end{cases}$
 $D^{\prime}$ $\coloneqq(A^{\prime}\sqcup B^{\prime},\operatorname{\lambda}^{\prime}).$
@@
 We give a general reduction rule which reduces the graph in case the instance is not extremally matched.
-This structural property, that we can assume the existence of such a perfect matching or reduce the instance, is useful for almost all further constructions.
+This structural property, that we can assume the existence of such perfect matchings or reduce the instance, is useful for almost all further constructions.
 For example, it enables us to reduce the instance size, when the number of vertices per side is not equal.
@@
 In the final solution for the whole bi-clique, vertex $a$ then has a temporal path to every vertex $b\in B$ via vertex $a^{\prime}$.
-Similarly, if a vertex $b\in B$ is dismountable, every temporal path that ends in a different vertex can be extended to be a temporal path ending in vertex $b$, and we can remove $b$ from the bi-clique.
+Similarly, if a vertex $b\in B$ is dismountable, every temporal path that ends in $b^{\prime}$ can be extended to be a temporal path ending in vertex $b$, and we can remove $b$ from the bi-clique.
 The following lemma is similar to the one in [6]. For completeness, we prove it again.
@@
 Due to Lemma 4.2 there must be a dismountable vertex in these cases.
-Since from now on, both partitions of $D$ have the same size after removing dismountable vertices, we write $n\coloneqq|A|=|B|$ and call $n$ the size of the instance.
-This justifies also that the definition of $\operatorname{\mathcal{D}}$ assumes equal partition sizes and gives the following corollary.
+Since from now on, both parts of $D$ have the same size after removing dismountable vertices, we write $n\coloneqq|A|=|B|$ and call $n$ the size of the instance.
+This justifies also that the definition of $\operatorname{\mathcal{D}}$ assumes equal part sizes and gives the following corollary.
 Corollary 4.6.
@@
 However, there are instances without this structure [6].
-Thus, we study an even more widely applicable structure, based on a what we call a pivot-edge.
-We exploit pivot-edges to apply the divide-and-conquer paradigm to find linear-size temporal spanners.
+Thus, we study an even more widely applicable structure, based on a what we call a $c$-pivot-edge or partial pivot-edge (when omitting the concrete parameter).
+We exploit partial pivot-edges to apply the divide-and-conquer paradigm to find linear-size temporal spanners.
 To the best of our knowledge, this paradigm has not yet been applied to this problem.
-As we see later, the absence of a pivot-edge yields strong structural properties and it heavily restricts the class of temporal cliques where no linear spanner can be found using current techniques.
-5.1 Pivot-Edges
+As we see later, the absence of a partial pivot-edge yields strong structural properties and it heavily restricts the class of temporal cliques where no linear spanner can be found using current techniques.
+5.1 Partial Pivot-Edges
 Recall the definition of $\operatorname{In}$ and $\operatorname{Out}$ in Section 2.
@@
 All vertices in $\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ can be organized in a tree $T$ rooted at $a$, s.t.
-$T$ only has edges with label at least $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$ and for each $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ there are temporal paths $a\leadsto v$ and $b\leadsto v$ in $G[T]$.
+$T$ only has edges with label at least $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$ and for each $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ there are temporal paths $a\leadsto v$ and $b\leadsto v$ in $T$.
 The analogous result holds for vertices in $\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$.
@@
 It was considered in [16] and lately has been used in [7], where foremost trees are applied to temporal cliques.
-We can directly use this idea to temporally connect all $u\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ to all $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using a linear number of edges.
+With this idea, we can temporally connect all $u\in\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)$ to all $v\in\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)$ using a linear number of edges.
 Corollary 5.4.
@@
 We name the set ${\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}$ the pivot-set of $e$.
-If $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$ for $c\in(0,1]$, we call $e$ a $c$-pivot-edge.
-Theorem 5.8 (Pivot Edge).
+If $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$ for $c\in(0,1]$, we call $e$ a $c$-pivot-edge.
+When omitting the concrete constant $c$, we refer to $e$ as a partial pivot-edge.
+Theorem 5.8.
 Let $c\in(0,1]$.
 Let $D=(A,B,\lambda)$ be an extremally matched bi-clique with $n\coloneqq|A|=|B|$.
-If there is a $c$-pivot-edge in $D$, we can split the instance into two sub-instances with overall number of vertices reduced by $cn$.
-For each removed vertex, we include at most $\frac{8}{c}$ edges.
-If applicable recursively, we get a spanner of size $\frac{16}{c}n$.
+If there is a $c$-pivot-edge in $D$, we can split the instance into two sub-instances with overall number of vertices reduced by $2cn$.
+Per removed vertex, we include at most $\frac{4}{c}$ edges.
+If applicable recursively, we get a spanner of size $\frac{8}{c}n$.
 Proof 5.9.
-Let $e\in{A}\otimes{B}$ be such that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq cn$.
-Note that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq\mathopen{}\mathclose{{}\left(2+c}\right)n$.
-By Theorem 5.6, the sub-instances have sides of overall size at most $2n-\frac{2+c}{2}n=\mathopen{}\mathclose{{}\left(1-\frac{c}{2}}\right)n$.
-This means that $cn$ vertices have been removed using at most $38n-$ 3 edges overall.
-For each such vertex, we include $\frac{8n-3}{cn}\leq\frac{8}{c}$ edges.
+Let $e\in{A}\otimes{B}$ be such that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$.
+Note that $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|=\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cup\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|+\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq\mathopen{}\mathclose{{}\left(2+2c}\right)n$.
+By Theorem 5.6, the sub-instances have sides of overall size at most $2n-\frac{2+2c}{2}n=\mathopen{}\mathclose{{}\left(1-c}\right)n$.
+This means that $2cn$ vertices have been removed using at most $38n-$ 3 edges overall.
+Per such vertex, we include $\frac{8n-3}{2cn}\leq\frac{4}{c}$ edges.
 To recursively apply Theorem 5.6, observe that we can reduce any dismountable vertex until both sub-instances are extremally matched.
-For each vertex removed in this way, we include $2\leq\frac{8}{c}$ edges.
-As the number of edges included per removed vertex in both ways does not depend on $n$, we include at most $\frac{8}{c}\mathopen{}\mathclose{{}\left|A\sqcup B}\right|=\frac{16}{c}n$ edges.
+For each vertex removed in this way, we include $2\leq\frac{4}{c}$ edges.
+As the number of edges included per removed vertex in both ways does not depend on $n$, we include at most $\frac{4}{c}\mathopen{}\mathclose{{}\left|A\sqcup B}\right|=\frac{8}{c}n$ edges.
 This theorem tells us that if we repeatedly find such a $c$-pivot-edge $e\in{A}\otimes{B}$ with $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|$ being at least a (previously fixed) $c$-fraction of the current size of the instance, we get a linear-size bi-spanner for our graph.
@@
 The following subsection gives an intuition for how restrictive this condition actually is.
-5.2 Graph Properties leading to Pivot-Edges
-Pivot-edges generalize pivotable vertices and yield a number of new, interesting structural properties in residual instances.
-In this section, we examine structures that, if present, imply that there is at least one pivot-edge.
-We first look at one particular case of pivot-edges, which we call steep edges.
+5.2 Graph Properties leading to Partial Pivot-Edges
+Partial pivot-edges generalize pivotable vertices and yield a number of new, interesting structural properties in residual instances.
+In this section, we examine structures that, if present, imply that there is at least one partial pivot-edge.
+We first look at one particular case of partial pivot-edges, which we call steep edges.
 Intuitively, an edge is steep, if it belongs to the later edges for one of the incident vertices and to the earlier edges for the other one.
@@
 Combining both inequalities, we get $\mathopen{}\mathclose{{}\left|\operatorname{In}\mathopen{}\mathclose{{}\left({e}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({e}}\right)}\right|\geq 2cn$.
-While steep edges demonstrate how an intuitively common structure cannot be there if assume there are no pivot-edges.
+While steep edges demonstrate how an intuitively common structure cannot be there if assume there are no partial pivot-edges.
 However, in randomly generated instances we expect edges not to be steep.
@@
 In fact, one can prove that with high probability, there is no $c$-steep in the graph.
-Yet, as every $c$-steep edge is pivotable, we want to analyze the graphs without such edges in greater detail.
+Yet, as every $c$-steep edge is a $c$-pivot-edge, we want to analyze the graphs without such edges in greater detail.
 We first look at more simple properties showing that any node only has a limited number of neighbors for which it is at least the $i$-th and at most the $j$-th neighbor.
@@
 6 Bi-Cliques With Reverted Edges
-Now that we have introduced the concept of $c$-pivot-edges and observed some interesting structural properties in bi-cliques which contain no suitable $c$-pivot-edge.
+We introduced partial pivot-edges and observed some interesting structural properties in bi-cliques which contain no suitable $c$-pivot-edge.
 The first hope would be that every bi-cliques admits such a $c$-pivot-edge.
-Unfortunately, this is not the case and we provide a general graph class, the shifted matching graph, in which for every edge $e$, the pivot-set is as small as possible (i.e., $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)=\{a,b\}$) and none of our reduction rules from Section 4 are applicable.
+Unfortunately, this is not the case and we provide a graph class, the shifted matching graph, in which for every edge, the pivot-set is as small as possible (i.e., $\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)\cap\operatorname{Out}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}}}\right)=\{a,b\}$) and none of our reduction rules from Section 4 apply.
 Still, we present a novel technique, called $e$-reverted edges, to construct small spanners for shifted matching graphs which also provides solutions for a significantly larger class of bi-cliques.
-We relate the technique to $c$-pivot-edges from Section 5 and argue why $e$-reverted edges solve a class distinct from $c$-pivot-edges, beyond the concrete example of shifted matching graphs.
+We relate the technique to partial pivot-edges from Section 5 and argue why $e$-reverted edges solve a class distinct from partial pivot-edges, beyond the concrete example of shifted matching graphs.
 Unfortunately, also this is not applicable to all graph classes, which we later demonstrate with a class that we call product graphs in Section 7.
@@
 Additionally, we observe that each edge at time $t$ has the form $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{i+t\operatorname{mod}n}}}\right\}$.
-Thus, the edges with label $t$ form a perfect matching and $a_{i}$ sees the $b_{j}$ in ascending order of indices, rotated by $i$ places compared to $a_{0}$.
+Thus, the edges with label $t$ form a perfect matching and $a_{i}$ sees the $b_{j}$ in ascending order of indices, circularly shifted by $i$ places compared to $a_{0}$.
 {observation}
 Let $t\in\{0,\dots n-1\}$ be a time-point. Then in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$, the edges with label $t$ form a perfect matching.
-We claim that this graph has no $c$-pivot-edges that are suitable for reduction in the sense of Theorem 5.8.
+We claim that this graph has no partial pivot-edges that are suitable for reduction in the sense of Theorem 5.8.
 To prove that this method fails on the shifted matching graph, we first characterize the set of vertices which can be reached from a vertex $a_{i}$ before and after a given time.
-Finally, we prove that there, the intersection of $\operatorname{In}$ and $\operatorname{Out}$ for all edges at $a_{i}$ is trivial.
+Finally, we prove that the intersection of $\operatorname{In}$ and $\operatorname{Out}$ for all edges at $a_{i}$ is trivial.
 Lemma 6.2.
@@
 No additional vertices can now reach $a_{i}$ and $b_{i+t+1\operatorname{mod}n}$, since after taking an edge with label $t+1$ from a different vertex, we are stuck due to Section 6.
-Using Lemma 6.2, we can conclude that using $c$-pivot-edges in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ leads to no reduction.
+Using Lemma 6.2, we can see that using partial pivot-edges in $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ leads to no reduction.
 Lemma 6.4.
@@
 By Lemma 6.2,
-$\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},\ldots,a_{i+t\operatorname{mod}n},b_{i\operatorname{mod}n},\ldots,b_{i+t\operatorname{mod}n}\}$
+$\operatorname{In}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}}}\right)=\{a_{i},\ldots,a_{i+t\operatorname{mod}n},b_{i},\ldots,b_{i+t\operatorname{mod}n}\}$
 and
@@
 […]
-Intuitively, for each edge $e$ there is a linear-size set of edges such that all pairs of vertices $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}$, that share an $e$-reverted edge, are connected via a short path.
+Intuitively, for each edge $e$ there is a linear-size set of edges that connects all pairs of vertices $\mathopen{}\mathclose{{}\left\{{a^{\prime}},{b^{\prime}}}\right\}$ that form an $e$-reverted edge.
 If ${a^{\prime}}\preceq_{b}{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)}$, the path $a^{\prime}\to b\to\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)\to b^{\prime}$ is temporal, since $b^{\prime}$ is also the latest neighbor for $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({b^{\prime}}}\right)$.
@@
 Figure 5: If at least one of the green or purple paths is temporal, $\{a^{\prime},b^{\prime}\}$ is $\mathopen{}\mathclose{{}\left\{{a},{b}}\right\}$-reverted.
-Notice that, no matter which $a^{\prime}\in A$ and $b^{\prime}\in B$ we choose, the paths have many edges in common.
-We always use edges incident to $a$ or $b$ and $\operatorname{\pi^{\text{--}}}$ or $\operatorname{\pi^{+}}$.
+Notice that, no matter which $a^{\prime}\in A$ and $b^{\prime}\in B$ we choose, the paths only use edges incident to $a$ or $b$ and edges from $\operatorname{\pi^{\text{--}}}$ or $\operatorname{\pi^{+}}$.
 We use this to construct a small bi-spanner for $D$, if $\operatorname{NotRev}_{{e}}$ is small.
@@
 We show the statement only for $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)=0$, the proof for label $n-1$ follows analogously.
-Let wlog.
-$a=a_{0}$ and $b=b_{0}$ and let $i,j\in\{0,\ldots,{n}-1\}$.
+Without loss of generality, let $a=a_{0}$ and $b=b_{0}$ and let $i,j\in\{0,\ldots,{n}-1\}$.
 To show that $\mathopen{}\mathclose{{}\left\{{a_{i}},{b_{j}}}\right\}$ is $e$-reverted, we distinguish two cases.
@@
 For every $n\in\operatorname{\mathbb{N}}^{+}$, the graph $\operatorname{SM}\mathopen{}\mathclose{{}\left({n}}\right)$ has a bi-spanner of size $44n-$ 4.
-Now we connect the concept of $e$-reverted edges to $c$-pivot-edges from Section 5
+Now we connect the concept of $e$-reverted edges to partial pivot-edges from Section 5.
 Lemma 6.12.
@@
 Lemma 6.12 tells us something about where to look for edges $e$ with many $e$-reverted edges.
-In fact, only edges that are either among the earliest or the latest edges of the incident vertices are suitable candidates for Theorem 6.7, which are the least likely to be $c$-pivot-edges.
+In fact, only edges that are either among the earliest or the latest edges of the incident vertices are suitable candidates for Theorem 6.7, which are the least likely to be suitable partial pivot-edges.
 This gives us an intuition, why the two techniques can be applied in different situations.
-This intuition is supported by the concrete example of shifted matching graphs.
+This intuition is supported by the example of shifted matching graphs.
 7 Product Graphs
@@
 Since $f(n)\in\operatorname{\mathcal{O}}(n)$, we have $s_{n}\in\Theta(n)$.
-By Lemma 7.8, we have that any pivot set is in $\operatorname{\mathcal{O}}(k_{n})=\operatorname{\mathcal{O}}(f(n))$, and by Lemma 7.10 any set of not reverted edges has size at least
+By Lemma 7.8, we have that any pivot-set is in $\operatorname{\mathcal{O}}(k_{n})=\operatorname{\mathcal{O}}(f(n))$, and by Lemma 7.10 any set of not reverted edges has size at least
 $(m_{n}-1)\binom{k_{n}}{2}$ $\geq\frac{n}{f(n)}\mathopen{}\mathclose{{}\left(\frac{(f(n)+1)\cdot f(n)}{2}}\right)\geq\frac{nf(n)}{2}.$
@@
 As $P_{H}$ is temporal, path $P$ remains temporal and it is easy to verify that $P$ is indeed a path from $(a_{G},a_{H})$ to $(b_{G},b_{H})$.
-[…]
+Case 2: $a_{G}^{\prime}\neq\pi^{+}_{{S_{G}}}({b_{G}})$.
+We append $\mathopen{}\mathclose{{}\left\{{(a_{G}^{\prime},a_{H})},{(b_{G},b_{H}^{*})}}\right\}$, which is in $S$ because $\mathopen{}\mathclose{{}\left\{{a_{G}^{\prime}},{b_{G}}}\right\}$ is the last edge of $P_{G}$.
+Further, we can append $\mathopen{}\mathclose{{}\left\{{(\pi^{+}_{{S_{G}}}({b_{G}}),a_{H}^{*})},{(b_{G},b_{H}^{*}}}\right\}$ to $P$, since $\mathopen{}\mathclose{{}\left\{{\pi^{+}_{{S_{G}}}({b_{G}})},{b_{G}}}\right\}\in S_{G}$.
+Again, as $S_{H}$ is a bi-spanner for $H$, there exists a temporal path $P^{\prime}_{H}$ from $a_{H}^{*}$ to $b_{H}$ in $S_{H}$.
+In the same fashion as in Case 1, for every edge $\mathopen{}\mathclose{{}\left\{{a_{H}^{P}},{b_{H}^{P}}}\right\}$ in $P^{\prime}_{H}$, we append $\mathopen{}\mathclose{{}\left\{{(\pi^{+}_{{S_{G}}}({b_{G}}),a_{H}^{P})},{(b_{G},b_{H}^{P})}}\right\}$ to $P$.
+See Figure 6(b) for a visualization of the constructed path $P$.
+As $\mathopen{}\mathclose{{}\left\{{a_{H}^{*}},{b_{H}^{*}}}\right\}$ has the minimal label in $S_{H}$ and $a_{G}^{\prime}\neq\pi^{+}_{{S_{G}}}({b_{G}})$, our construction yields a temporal path which is clearly included in $S$.
 If $G$ and $H$ each have linear-size bi-spanners, this proves the existence of a linear-size bi-spanner for ${G}\operatorname{\times}{H}$.
@@
 The red area contains instances that are not solved by any known techniques.
-We have made a significant step forwards towards answering whether temporal cliques admit linear-size temporal spanners, see Figure 7 for an illustration.
+We present a significant step forwards towards answering whether temporal cliques admit linear-size temporal spanners, see Figure 7 for an illustration.
 The obvious problem that stems from this paper, is whether classes of graphs exist that are not solved by our techniques.
@@
 Future research could generalize this idea and more closely analyze the cases in which a temporal spanner can be obtained by combining temporal spanners on subinstances.
-Towards the opposite direction, it is now easier to find a $\Omega(n\log n)$ counterexample, if it exists at all, since we have provided many properties that cannot exist in a counterexample.
+Towards the opposite direction, it is now easier to find a $\Omega(n\log n)$ counterexample, if it exists at all, since we have provided many properties that any counterexample must avoid.